### v1

# DAISI: Data Assimilation with Inverse Sampling using Stochastic Interpolants

Abstract
Data assimilation (DA) is a cornerstone of scientific and engineering applications, combining model forecasts with sparse and noisy observations to estimate latent system states. Classical DA methods, such as the ensemble Kalman filter, rely on Gaussian approximations and heuristic tuning (e.g., inflation and localization) to scale to high dimensions. While often successful, these approximations can make the methods unstable or inaccurate when the underlying distributions of states and observations depart significantly from Gaussianity. To address this limitation, we introduce DAISI, a scalable filtering algorithm built on flow-based generative models that enables flexible probabilistic inference using data-driven priors. The core idea is to use a stationary, pre-trained generative prior to assimilate observations via guidance-based conditional sampling while incorporating forecast information through a novel inverse-sampling step. This step maps the forecast ensemble into a latent space to provide initial conditions for the conditional sampling, allowing us to encode model dynamics into the DA pipeline without having to retrain or fine-tune the generative prior at each assimilation step. Experiments on challenging nonlinear systems show that DAISI achieves accurate filtering results in regimes with sparse, noisy, and nonlinear observations where traditional methods struggle.
Machine Learning, ICML
1 Introduction
Figure 1: Our method, DAISI, performs data assimilation by combining a flow-based unconditional prior $\mathbb{P}_{\infty}$ with the forecast ensemble $\hat{\pi}_{n}$. To condition on an observation ${\bm{y}}_{n}$, one could apply the guided SDE starting from random noise, producing $\mathbb{P}_{\infty}^{{\bm{y}}}$ and ignoring the forecast (blue). Instead, DAISI (green): (i) applies the backward SDE to the forecast ensemble, producing inverted samples $\Psi_{\sharp}\hat{\pi}_{n}$, and (ii) uses these latents as initial conditions for the guided SDE, generating approximate samples from the filtering distribution $\pi_{n}$.
Estimating the evolving state of a complex dynamical system is a fundamental challenge in science and engineering. In many real-world problems—such as weather forecasting, fluid mechanics, neuroscience and robotics—the states of interest are only partially observed and subject to multiple sources of uncertainty (asch_DA_2016). DA, and in particular, filtering, seeks to combine imperfect model forecasts with sparse, noisy observations to reconstruct these hidden states. Mathematically, this corresponds to estimating the filtering distribution $p({\bm{x}}_{n}|{\bm{y}}_{1:n})$ where ${\bm{x}}_{n}$ denotes the latent state and ${\bm{y}}_{1:n}:=({\bm{y}}_{1},\,\dots,\,{\bm{y}}_{n})$, where ${\bm{y}}_{n}$ is the observation at time $n$ generated via the likelihood $p({\bm{y}}_{n}|{\bm{x}}_{n})$.
Weather forecasting provides a clear illustration of the challenges present in DA. The atmosphere is chaotic, high-dimensional, and observed through nonlinear, noisy measurements. Accurate state estimation is essential for applications ranging from early warning systems and agriculture to aviation safety and renewable energy management. (noaa2025billiondollar; whitt2023economic; ipcc2023synthesis; energy_pred). To meet these demands, classical DA methods, such as the Ensemble Kalman Filter (EnKF), or variational methods such as 4DVar, have been developed over decades of DA research. However, these have respective drawbacks; EnKF is only guaranteed to work under near-Gaussian settings (calvello2024accuracy), and the inflation and localization parameters necessary to stabilize the filter are notoriously challenging to tune (bannister2017review). 4DVar, on the other hand, requires the tedious development of an adjoint model and cannot quantify uncertainty, due to it being a MAP estimation method. Particle filters (naesseth2019elements) can, in principle, solve the filtering problem, but suffers from the curse of dimensionality (bengtsson2008curse). These limitations motivate the development of more flexible, data-driven approaches to DA.
Recent progress in offline inverse problems—such as those in medical imaging, astrophysics, and computer vision—has shown that flow- and diffusion-based generative models can serve as powerful priors for conditional sampling (zhenginversebench). However, extending these ideas to sequential filtering presents unique challenges. One possibility is to learn a generative prior for the predictive distribution $p({\bm{x}}_{n}|{\bm{y}}_{1:n-1})$, but this would require retraining the model at every time step $n$ (bao_1_trained) , making it impractical for operational use. Another strategy is to use a pre-trained generative forecast model to guide towards observations (chen2025flowdas; savary2025training), though such models require the use of specialized dynamical models. Alternatively, approaches based on approximating the smoothing distribution $p({\bm{x}}_{0:n}|{\bm{y}}_{1:n})$ (rozet2023score) can, in principle, be used for filtering. However, they are memory-intensive and scale poorly with the number of time steps. We elaborate on how these approaches relate to our proposed method in Section 3.3.
1.1 Contributions
In this work, we present DAISI (Data Assimilation with Inverse sampling using Stochastic Interpolants), a scalable filtering algorithm built on flow-based generative models. DAISI avoids retraining a generative model at each assimilation step by leveraging a stationary pre-trained generative prior to condition on new observations via guidance. To integrate dynamical information, DAISI couples this generative prior with a forecast model that advances an ensemble of states in time. Following the forecast, an inverse-sampling step runs the generative SDE backward from the forecast ensemble, mapping forecasted states to latent variables. These latent states then serve as initial conditions for conditional sampling under the learned prior. The full procedure is illustrated in Figure 1.
In summary, the strength and novelty of DAISI lie in its following capabilities:
  1. (i)
    Zero-shot compatibility with both numerical and ML-based forecast and observation models.
  2. (ii)
    Modular design, supporting any flow-based generative model or gradient-based guidance method.
  3. (iii)
    Expressive uncertainty quantification, capturing complex, multimodal posteriors under sparse, noisy, and nonlinear observations.
2 Preliminaries
2.1 Data Assimilation
Consider the state-space model
${\bm{x}}_{n}$ $=\mathcal{F}({\bm{x}}_{n-1},\bm{\omega}_{n}),\quad\bm{\omega}_{n}\sim p(\bm{\omega})$ (1)
${\bm{y}}_{n}$ $=\mathcal{H}({\bm{x}}_{n})+\bm{\nu}_{n},\quad\bm{\nu}_{n}\sim p(\bm{\nu})$ (2)
defined for $n=1,\ldots,N$ and ${\bm{x}}_{0}\sim p({\bm{x}}_{0})$. Here, $\mathcal{F}:\mathcal{X}\times{\Omega}\rightarrow\mathcal{X}$ and $\mathcal{H}:\mathcal{X}\rightarrow\mathcal{Y}$ are the dynamics and observation operators, respectively, and $\bm{\omega}_{n}\in{\Omega},\bm{\nu}_{n}\in\mathcal{Y}$ are stochastic noise sources. Our general formulation of the state evolution (1) encompasses various settings such as a deterministic physics-based simulator (in which case $\bm{\omega}_{n}\equiv 0$) or a pre-trained generative model (in which case $\bm{\omega}_{n}$ corresponds to the driving noise process of the model).
DA seeks to infer the latent states $\{{\bm{x}}_{n}\}_{n=0}^{N}$ from the noisy observations $\{{\bm{y}}_{n}\}_{n=1}^{N}$. In particular, filtering refers to the problem of estimating the distribution $p({\bm{x}}_{n}|{\bm{y}}_{1:n})$, while smoothing targets the full posterior $p({\bm{x}}_{0:N}|{\bm{y}}_{1:N})$. In this work, we focus on the filtering problem, generally solved by alternating between a forecast and analysis step. Denoting by $\pi_{n}(\mathrm{d}{\bm{x}}_{n}):=p({\bm{x}}_{n}|{\bm{y}}_{1:n})\mathrm{d}{\bm{x}}_{n}$, the measure corresponding to the filtering distribution, this proceeds abstractly as
$\text{(Forecast)}\,\,\,\,\hat{\pi}_{n}(\mathrm{d}{\bm{x}}_{n}):=\mathbb{E}_{\bm{\omega}_{n}}\left[\mathcal{F}(\cdot,\bm{\omega}_{n})_{\sharp}\pi_{n-1}\right](\mathrm{d}{\bm{x}}_{n})$ (3)
$\text{(Analysis)}\,\,\,\,\pi_{n}(\mathrm{d}{\bm{x}}_{n})\propto p({\bm{y}}_{n}|{\bm{x}}_{n})\hat{\pi}_{n}(\mathrm{d}{\bm{x}}_{n}),$ (4)
where $\mathcal{F}_{\sharp}\pi$ denotes the pushforward of the measure $\pi$ with respect to a map $\mathcal{F}$, defined by $\mathcal{F}_{\sharp}\pi:=\mathrm{Law}(\mathcal{F}(X))$ for $X\in\pi$. Starting from $\pi_{0}$, repeated applications of (3)–(4) yields the sequence $\pi_{1},\ldots,\pi_{N}$. Although this recursion is well defined, computing these measures exactly is typically infeasible in practice. Our aim is therefore to develop a practical and scalable approximation to the cycle (3)–(4), capable of operating reliably in high-dimensional settings with sparse, nonlinear observations.
2.2 Stochastic Interpolants for Generative Modeling
Our proposed DA method leverages recent advances in generative modeling. In the static setting (i.e., ignoring time evolution), flow-based models enable sampling from a target distribution $\rho_{1}$ by learning a transport map from a simple latent distribution $\rho_{0}$. Such transports can be parameterized in multiple ways; in this work, we adopt the framework of stochastic interpolants (albergo2023stochastic), though closely related formulations appear in the probability flow ODE for diffusion models (song2020score) and in flow matching (lipmanflow).
We define a stochastic interpolant as a stochastic process
${\bm{z}}_{t}=\alpha_{t}{\bm{z}}_{1}+\beta_{t}{\bm{z}}_{0},\quad t\in[0,1],$ (5)
where the curves $\alpha,\beta\in C^{2}([0,1])$ satisfy the endpoint conditions $\alpha_{0}=\beta_{1}=0$ and $\alpha_{1}=\beta_{0}=1$. The data pair $({\bm{z}}_{0},{\bm{z}}_{1})$ is sampled from a measure $\nu(\mathrm{d}{\bm{z}}_{0},\mathrm{d}{\bm{z}}_{1})$ whose marginals correspond to $\rho_{0}(\mathrm{d}{\bm{z}}_{0})$ and $\rho_{1}(\mathrm{d}{\bm{z}}_{1})$, respectively.
Let $\rho_{t}:=\mathrm{Law}({\bm{z}}_{t})$ denote the distribution of ${\bm{z}}_{t}$ with density $p_{t}$, and define its score ${\bm{s}}(t,{\bm{z}}):=\nabla\log p_{t}({\bm{z}})$ and drift ${\bm{b}}(t,{\bm{z}}):=\mathbb{E}[\dot{{\bm{z}}}_{t}|{\bm{z}}_{t}={\bm{z}}]$. In albergo2023stochastic, it is shown that for any non-negative $\epsilon\in C^{0}([0,1])$, the SDE
$\mathrm{d}{\bm{z}}_{t}=({\bm{b}}(t,{\bm{z}}_{t})+\epsilon_{t}{\bm{s}}(t,{\bm{z}}_{t}))\mathrm{d}t+\sqrt{2\epsilon_{t}}\mathrm{d}W_{t},$ (6a)
${\bm{z}}_{0}\sim\rho_{0},\quad t\in[0,1],$ (6b)
where $W_{t}$ is the Brownian motion, has the same marginal laws $\rho_{t}$ as the interpolant in (5). Furthermore, this is also true for the corresponding reverse-time process,
$\mathrm{d}{\bm{z}}_{t}=({\bm{b}}(t,{\bm{z}}_{t})-\epsilon_{t}{\bm{s}}(t,{\bm{z}}_{t}))\mathrm{d}t+\sqrt{2\epsilon_{t}}\mathrm{d}\widehat{W}_{t},$ (7a)
${\bm{z}}_{1}\sim\rho_{1},\quad t\in[0,1],$ (7b)
where $\widehat{W}_{t}$ is the Brownian motion in reverse time. In the important special case when $\rho_{0}=\mathcal{N}(\bm{0},\mathbf{I})$, the score ${\bm{s}}(t,{\bm{z}})$ can be explicitly related to the drift ${\bm{b}}(t,{\bm{z}})$. This enables sampling via (6) and (7) for arbitrary $\epsilon_{t}$ using only the drift, which can be learned from samples $({\bm{z}}_{0},{\bm{z}}_{1})\sim\nu$; See Appendix A.1 for details.
2.3 Conditional Generation via Guidance
Given an observation ${\bm{y}}\sim p({\bm{y}}|{\bm{x}})$ and a forward SDE (6), we can sample from the posterior measure $\rho^{{\bm{y}}}_{1}(\mathrm{d}{\bm{x}})\propto p({\bm{y}}|{\bm{x}})\rho_{1}(\mathrm{d}{\bm{x}})$ by solving the following forward SDE with guidance:
$\mathrm{d}{\bm{z}}_{t}=(\tilde{{\bm{b}}}(t,{\bm{z}}_{t};{\bm{y}})+\epsilon_{t}\tilde{{\bm{s}}}(t,{\bm{z}}_{t};{\bm{y}}))\mathrm{d}t+\sqrt{2\epsilon_{t}}\mathrm{d}W_{t},$ (8a)
${\bm{z}}_{0}\sim\rho_{0},\quad t\in[0,1],$ (8b)
where the guided score $\tilde{{\bm{s}}}$ is given by
$\tilde{{\bm{s}}}(t,{\bm{z}}_{t};{\bm{y}})={\bm{s}}(t,{\bm{z}}_{t})+\nabla_{{\bm{z}}_{t}}\log p({\bm{y}}|{\bm{z}}_{t}),$ (9)
and the guided drift $\tilde{{\bm{b}}}$ by
$\tilde{{\bm{b}}}(t,{\bm{z}}_{t};{\bm{y}})={\bm{b}}(t,{\bm{z}}_{t})+\lambda_{t}\nabla_{{\bm{z}}_{t}}\log p({\bm{y}}|{\bm{z}}_{t}),$ (10)
for $\lambda_{t}$ defined as in Appendix A.4.
In general, the guidance term is intractable as we usually only have access to $p({\bm{y}}|{\bm{z}}_{1})$ but not $p({\bm{y}}|{\bm{z}}_{t})=\mathbb{E}_{{\bm{z}}_{1}|{\bm{z}}_{t}}[p({\bm{y}}|{\bm{z}}_{1})]$. Various approaches exist to approximate this term, e.g. diffusion posterior sampling (DPS) (chung2023diffusion) and moment matching posterior sampling (MMPS) (Rozet_Andry_Lanusse_Louppe_2024) (see Appendix A.5 for details and (daras2024survey) for a survey on guidance methods).
3 Method
Our method, DAISI, performs filtering by combining the strengths of ensemble-based methods with the expressive priors offered by flow-based generative models. This is conceptually similar to the hybrid ensemble-variational (EnVar) method in classical DA (hamill2000hybrid; lorenc2003potential), which blend a static climatological background covariance $\mathbf{P}$ with an “error-of-the-day” ensemble covariance to represent forecast uncertainty. This compensates for the limitations of either component alone: relying solely on a static background covariance ignores the dynamical evolution of errors, while using only a small ensemble fails to characterize high-dimensional uncertainties accurately.
DAISI builds on this philosophy but moves beyond Gaussian background covariances $\mathbf{P}$ used in hybrid EnVar by considering a full background measure $\mathbb{P}_{\infty}$, learned using flow-based generative models. In practice, we take $\mathbb{P}_{\infty}$ to be the invariant measure of the dynamical system (1), whose samples can be generated via long-time rollouts. While its existence and ergodicity are not generally guaranteed, it is a reasonable assumption for many geophysical flows, which are typically dissipative and subject to stochastic forcing. Given the learned prior $\mathbb{P}_{\infty}$, we can then assimilate new observations ${\bm{y}}_{n}$ by sampling from the posterior
$\mathbb{P}_{\infty}^{\bm{y}}(\mathrm{d}{\bm{x}}_{n})\propto p({\bm{y}}_{n}|{\bm{x}}_{n})\mathbb{P}_{\infty}(\mathrm{d}{\bm{x}}_{n})$ (11)
using guidance (Section 2.3). However, this alone neglects the dynamical evolution encoded in (1), which is essential for accurately tracking the latent states through time. This necessitates coupling $\mathbb{P}_{\infty}$ with the forecast ensemble, echoing the role of the ensembles in hybrid EnVar schemes.
To achieve this, we introduce inverse sampling, which transfers dynamical information from the ensemble forecast into the latent space of the generative model. Given forecast particles $\{\hat{{\bm{x}}}^{(j)}_{n}\}_{j=1}^{J}$, this proceeds by solving the unconditional backward SDE (7) from $t=1$ to $t=t_{\min}\in(0,1)$, using each $\hat{{\bm{x}}}_{n}^{(j)}$ as terminal conditions. The resulting latent variables $\{{\bm{z}}_{t_{\min},n}^{(j)}\}_{j=1}^{J}$ encode the forecast information in “noise space”. We can then perform conditional sampling by integrating the guided forward SDE (8) from $t_{\min}$ to 1, initialized at these latent states. This produces updated particles $\{{\bm{x}}_{n}^{(j)}\}_{j=1}^{N}$ that are samples of $\mathbb{P}_{\infty}^{\bm{y}}$, while simultaneously containing information about the ensemble $\{\hat{{\bm{x}}}^{(j)}_{n}\}_{j=1}^{J}$ through its latent representations $\{{\bm{z}}_{t_{\min},n}^{(j)}\}_{j=1}^{J}$.
Taken altogether, DAISI performs filtering by alternating between the following forecast and analysis steps:
Forecast.
Given particles $\{{\bm{x}}^{(j)}_{n-1}\}_{j=1}^{J}$ approximating the filtering distribution ${\pi}_{n-1}$, generate forecasts
$\hat{{\bm{x}}}_{n}^{(j)}$ $=\mathcal{F}({\bm{x}}_{n-1}^{(j)},\bm{\omega}_{n}^{(j)}),\quad j=1,\ldots,J,$ (12)
for i.i.d. noise realizations $\bm{\omega}_{n}^{(j)}\sim p(\bm{\omega})$. This produces samples from the predictive distribution $\hat{\pi}_{n}$ (see Equation 3).
Analysis.
Assuming access to a flow-based generative model bridging $\rho_{0}=\mathcal{N}(\bm{0},\mathbf{I})$ to $\rho_{1}=\mathbb{P}_{\infty}$, and given forecast particles $\{\hat{{\bm{x}}}_{n}^{(j)}\}_{j=1}^{J}$:
  1. 1.
    Inverse sampling: Solve the backward SDE (7) from $t=1$ to $t_{\min}$ with terminal conditions $\hat{{\bm{x}}}_{n}^{(j)}$, to obtain the latent states ${\bm{z}}_{t_{\min},n}^{(j)}$.
  2. 2.
    Forward guided sampling: Solve the conditional forward SDE (8), from $t=t_{\min}$ to 11 with initial condition ${\bm{z}}^{(j)}_{t_{\min},n}$ to obtain updated particles ${\bm{x}}_{n}^{(j)}$.
Algorithm 1 DAISI filtering
1:In: Samples $\{{\bm{x}}_{0}^{(j)}\}_{j=1}^{J}\sim\pi_{0}$, observations $\{{\bm{y}}_{n}\}_{n=1}^{N}$
2:for $n=1,\dots,N$ do
3:  Forecast:
4:  for $j=1,\dots,J$ do
5:   Sample $\bm{\omega}_{n}^{(j)}\sim p(\bm{\omega})$
6:   $\widehat{{\bm{x}}}_{n}^{(j)}\leftarrow\mathcal{F}({\bm{x}}_{n-1}^{(j)},\bm{\omega}_{n}^{(j)})$
7:  end for
8:  Inverse sampling:
9:  for $j=1,\dots,J$ do
10:   Integrate SDE (7) backward from $t=1$ to $t_{\min}$ with terminal state $\widehat{{\bm{x}}}_{n}^{(j)}$ to obtain ${\bm{z}}^{(j)}_{t_{\min},n}$
11:  end for
12:  Guided sampling:
13:  for $j=1,\dots,J$ do
14:   Integrate guided SDE (8) from $t_{\min}$ to 11 with initial state ${\bm{z}}^{(j)}_{t_{\min},n}$ and observation ${\bm{y}}_{n}$ to obtain ${\bm{x}}_{n}^{(j)}$
15:  end for
16:end for
17:Out: Updated particles $\{{\bm{x}}_{n}^{(j)}\}_{j=1}^{J}$ for $n=1,\ldots,N$
We summarize this assimilation procedure in Algorithm 1.
3.1 Does DAISI sample from the filtering distribution?
For $0<s\leq t\leq 1$ and $\epsilon\geq 0$, denote by $\Phi_{s,t}^{\epsilon},\Phi_{s,t}^{{\bm{y}},\epsilon}$ the stochastic flows of the forward SDEs (6), (8), respectively. Likewise, let $\Psi_{t,s}^{\epsilon},\Psi_{t,s}^{{\bm{y}},\epsilon}$ denote the corresponding backward-flow maps. We assume for simplicity that the noise coefficient in the SDEs are given by constants $\epsilon_{t}\equiv\epsilon$. For a fixed time step $n$ and some $t_{\min}\in(0,1)$, we see that the analysis step in DAISI samples from the measure
$\pi_{n,t_{\min},\epsilon}^{\text{DAISI}}:=\mathbb{E}\left[(\Phi_{t_{\min},1}^{{\bm{y}},\epsilon}\circ\Psi_{1,t_{\min}}^{\epsilon})_{\sharp}\hat{\pi}_{n}\right],$ (13)
where $\hat{\pi}_{n}$ denotes the predictive distribution at time $n$, and the expectation is over the Brownian motions used when integrating the backward and forward SDEs. For DAISI to function as a reliable filter, this measure should approximate the true filtering distribution $\pi_{n}$ closely. While there is no reason to believe that (13) matches $\pi_{n}$ exactly, we argue that we can achieve a reasonably good match by tuning the hyperparameters $t_{\min}$ and $\epsilon$ appropriately.
In the following, let us assume that for all time $n$, the predictive distribution $\hat{\pi}_{n}$ is absolutely continuous with respect to the invariant measure $\mathbb{P}_{\infty}$, i.e., there exists a measurable density ratio $f_{n}:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}_{\geq 0}$ such that $\hat{\pi}_{n}(\mathrm{d}{\bm{x}}_{n})\propto f_{n}({\bm{x}}_{n})\mathbb{P}_{\infty}(\mathrm{d}{\bm{x}}_{n})$. Then, we can rewrite the filtering distribution as follows
[…]
$\propto p({\bm{y}}_{n}|{\bm{x}}_{n})f_{n}({\bm{x}}_{n})\mathbb{P}_{\infty}(\mathrm{d}{\bm{x}}_{n})$ (15)
$\propto f_{n}({\bm{x}}_{n})\mathbb{P}^{\bm{y}}_{\infty}(\mathrm{d}{\bm{x}}_{n}).$ (16)
We now compare the DAISI analysis distribution (13) with the ideal target (16) in various limits to understand the role of each hyperparameter.
Effect of $t_{\min}$.
We first examine the limiting cases $t_{\min}\rightarrow 1$ and $t_{\min}\rightarrow 0$, assuming $\epsilon=0$ (deterministic flows). By time-continuity of the flows on $[0,1]$, the first limit $t_{\min}\rightarrow 1$ trivially yields $\pi_{n,t_{\min},0}^{\text{DAISI}}\rightarrow\hat{\pi}_{n}\propto f_{n}\mathbb{P}_{\infty}$. Comparing with (16), we observe that although the density ratios agree, the underlying base measures differ.
Now for the limit $t_{\min}\rightarrow 0$, first note that the inverse sampling step in DAISI produces the intermediary measure
$\widetilde{\rho}_{0}$ $:=(\Psi_{1,0}^{0})_{\sharp}\hat{\pi}_{n}\propto(f_{n}\circ(\Psi_{1,0}^{0})^{-1})(\Psi_{1,0}^{0})_{\sharp}\mathbb{P}_{\infty}$
$=(f_{n}\circ(\Psi_{1,0}^{0})^{-1})\rho_{0},$ (17)
where we used that $(\Psi_{1,0}^{0})_{\sharp}\mathbb{P}_{\infty}=\rho_{0}$. Next applying the guided flow in the second step gives us
$\pi_{n,0,0}^{\text{DAISI}}=(\Phi^{{\bm{y}},0}_{0,1})_{\sharp}\widetilde{\rho}_{0}$ (18)
[…]
where $g_{n}^{{\bm{y}}}:=f_{n}\circ(\Phi^{{\bm{y}},0}_{0,1}\circ\Psi^{0}_{1,0})^{-1}$ and we used that $\Phi^{{\bm{y}},0}_{0,1}$ transports $\rho_{0}$ to $\mathbb{P}_{\infty}^{\bm{y}}$ by construction. Comparing (19) with (16), we see that in this limit, the base measures match, while the density ratios differ. Thus, the extremes $t_{\min}\rightarrow 1$ and $t_{\min}\rightarrow 0$ each match only one component of the true filtering distribution, suggesting the existence of an intermediate $t_{\min}\in(0,1)$ that provides the best trade-off between these two types of mismatch.
In Figure 2, we illustrate this trade-off using a simple 1D toy problem (see Appendix D.2 for details). When $t_{\min}=0.01$, the distirbution $\pi^{\text{DAISI}}_{n}$ (in orange) shares characteristics of $\mathbb{P}^{\bm{y}}_{\infty}$, as predicted from (19). Conversely, for a large value such as $t_{\min}=0.6$, the distribution closely matches $\hat{\pi}_{n}$, as expected. The intermediate setting $t_{\min}=0.3$ yields a distribution that interpolates between these two extremes, yielding a distribution that better aligns with the filtering distribution, which combines features of both $\mathbb{P}^{\bm{y}}_{\infty}$ and $\hat{\pi}_{n}$ — as reflected by the noticeably lower Maximum Mean Discrepancy (MMD) score (gretton2012kernel).
(a) $\quad t_{\min}=0.01$
(b) $\quad t_{\min}=0.3$
(c) $\quad t_{\min}=0.6$
Figure 2: Ablation with respect to $t_{\min}$, fixing $\epsilon=0$. The measure $\pi^{\text{DAISI}}_{n,t_{\min},\epsilon}$ (orange) is pulled towards $\hat{\pi}_{n}$ (green) as $t_{\min}\rightarrow 1$. There is an intermediate $t_{\min}^{*}$ where it matches $\pi_{n}$ (blue) the best.
Effect of $\epsilon$.
We next examine the effect of the noise parameter $\epsilon$. In the absence of observations, applying the backward process (7) to the forecast samples, followed by the forward process (6), provides a mechanism to resample the forecast ensemble. In the deterministic case $\epsilon=0$, the original ensemble members are exactly recovered. Introducing a small amount of noise via $\epsilon$ allows the generation of “new” forecast samples that are likely under $\mathbb{P}_{\infty}$, yet remains close to the original forecasts. As $\epsilon$} increases, the resampled ensemble progressively deviates, and in the limit $\epsilon$$\to\infty$, all ensemble information is lost. In the conditional setting, $\epsilon$ therefore controls the extent to which conditional samples retain forecast information. This can be stated more precisely in the following result.
Proposition 3.1.
Assume that the marginal laws $\{\rho_{t}\}_{t\in[0,1]}$ and $\{\rho_{t}^{\bm{y}}\}_{t\in[0,1]}$ of the interpolants bridging $\rho_{0}=\mathcal{N}(\bm{0},\mathbf{I})$ with $\rho_{1}=\mathbb{P}_{\infty}$ and $\rho_{1}=\mathbb{P}_{\infty}^{\bm{y}}$, respectively, satisfy a log-Sobolev inequality with constant $\lambda>0$, uniform in $t$. Then,
$\mathcal{KL}\left(\pi_{n,0,\epsilon}^{\text{DAISI}}\,||\,\mathbb{P}_{\infty}^{\bm{y}}\right)\leq e^{-4\lambda\epsilon}\,\mathcal{KL}\left(\hat{\pi}_{n}\,||\,\mathbb{P}_{\infty}\right),$ (20)
with equality at $\epsilon=0$, where $\mathcal{KL}(q||p)$ denotes the Kullback-Leibler divergence between measures $q$ and $p$.
Proof.
See Appendix C. ∎
We see from this result that in the case $\epsilon=0$, the Kullback–Leibler divergence between the DAISI output and the guided reference measure is preserved, so the conditioning step retains all information carried by the forecast ensemble. On the other hand, when $\epsilon\rightarrow\infty$, we have $\pi_{n,0,\epsilon}^{\text{DAISI}}\rightarrow\mathbb{P}_{\infty}^{\bm{y}}$ and therefore all ensemble information is lost. In practice, we find that taking a small but non-zero $\epsilon$ is useful for mitigating the tendency of particles from collapsing to a single state as assimilation progresses. In our 1D toy experiment (Appendix D.2), we show that tuning both $t_{\min}$ and $\epsilon$ yields better performance than adjusting $t_{\min}$ alone.
3.2 Complexity Analysis
Assuming the dynamics model $\mathcal{F}(\cdot)$ incurs a cost of $\mathcal{O}(f(d_{x}))$, the prediction step (12) costs $\mathcal{O}(Jf(d_{x}))$. Denoting by $\mathcal{O}(g(d_{x},d_{y}))$ the cost of a single Euler–Maruyama (EM) step for solving (6) or (7) (with or without guidance), the analysis step costs $\mathcal{O}(JTg(d_{x},d_{y}))$, where $T$ is the number of EM steps. Hence, the overall computational complexity of DAISI is $\mathcal{O}\big(NJ(f(d_{x})+Tg(d_{x},d_{y}))\big)$. For comparison, the ensemble transform Kalman filter (ETKF) has cost $\mathcal{O}\big(NJ(f(d_{x})+(d_{x}+d_{y})J+J^{2})\big)$. Since the forward evaluation of the U-net parameterizing the drift ${\bm{b}}(t,{\bm{z}})$ scales linearly with input dimension (i.e., $g(d_{x},d_{y})=\mathcal{O}(d_{x}+d_{y})$), DAISI achieves comparable cost to ETKF when $T\sim J$, while scaling linearly in ensemble size JJ—in contrast to ETKF’s cubic scaling in $J$. However, in settings where $J$ is small (e.g., $\mathcal{O}(10)$), DAISI can be more expensive due to the need for $T=\mathcal{O}(10^{2})$ EM steps to accurately solve the forward and backward SDEs.
We note that the Forecast, Inverse sampling, and Guided sampling steps in Algorithm 1 are all embarrassingly parallelizable, enabling further reductions in computational cost at the expense of increased memory use.
3.3 Related Works
Early work on diffusion priors for data assimilation includes Score-based Data Assimilation (SDA) (rozet2023score), which learns the score of full trajectories to perform smoothing without an explicit forecast model. While effective, it is memory-intensive and scales poorly for filtering.
Closer to our setting, Yang et al. (yang2025generative) address linear filtering by guiding a pre-trained diffusion model using both forecasts and observations. Instead of solving the backward SDE, they rely on SDEdit (meng2022sdedit), which partially inverts a sample by re-noising it, and incorporate observations via RePaint (lugmayr2022repaint), an iterative noising-denoising procedure. Together, these steps create a trade-off between preserving forecast information and enforcing the conditioning. In Appendix D.4.6, we show that replacing the backward SDE with SDEdit leads to substantially worse results.
The Ensemble Score Filter (EnSF) (bao2024ensemble) replaces the learned score with an analytical approximation, enabling training-free assimilation of nonlinear observations. However, it struggles in sparse settings due to the lack of a learned structure. An extension of EnSF in latent-space (si2025latentensf) partially addresses this; however, this does not easily amortize over observation models and its results depend heavily on the quality of the trained autoencoder. EnSF and its variants have been applied to the SQG model (bao2025nonlinear; yin2024scalable; liang2025ensemble), and extensions with Gaussian mixture score approximations (zhang2025iensf) and problem-dependent data coupling (transue2025flow) have been proposed.
Finally, recent works explore guidance for diffusion-based forecasting models. For instance, FlowDAS (chen2025flowdas) employs stochastic interpolants, while (savary2025training) uses guidance and particle filters to steer the GenCast model (gencast). These methods depend heavily on the accuracy of the learned one-step forecast distribution $p({\bm{x}}_{n+1}|{\bm{x}}_{n})$. When this distribution is too concentrated, the model effectively “expects” the next state to lie within a narrow range, and struggles to correct toward observations that fall outside it. Because DAISI learns the marginal distribution instead, it is less prone to such failures.
4 Experiments
We evaluate DAISI on three systems: the Lorenz ’63 (L63) system (lorenz1963deterministic), a Surface Quasi-Geostrophic system (SQG) (tulloch2009quasigeostrophic), and SEVIR, a real-world radar dataset (veillette2020sevir).
4.1 Lorenz ’63 System
RMSE ($\downarrow$) CRPS ($\downarrow$) SSR ($\approx 1$)
BPF $\mathbf{1.18}_{\pm 0.50}$ $\mathbf{1.46}_{\pm 0.49}$ $2.57_{\pm 0.56}$
DAISI (no inversion) $5.90_{\pm 1.09}$ $7.17_{\pm 1.38}$ $\underline{1.40}_{\pm 0.25}$
DAISI ($t_{\min}\approx 0$, $\epsilon=0$) $7.04_{\pm 0.90}$ $11.9_{\pm 1.55}$ $0.06_{\pm 0.02}$
DAISI (tuned $t_{\min}$, $\epsilon=0$) $2.44_{\pm 2.89}$ $3.58_{\pm 4.79}$ $\mathbf{1.07}_{\pm 1.09}$
DAISI (tuned both $t_{\min}$ & $\epsilon$) $\underline{2.03}_{\pm 1.01}$ $\underline{2.61}_{\pm 1.38}$ $\mathbf{1.08}_{\pm 0.41}$
DAISI (tuned both, set $\epsilon=0$) $2.05_{\pm 1.31}$ $3.19_{\pm 2.14}$ $0.32_{\pm 0.09}$
Table 1: Summary metrics on the L63 example. We display the mean and standard deviation across 10 independent experiments, over the last $100$ assimilation steps. The best score for each metric is highlighted in bold and the second best with an underline.
(a)   BPF
(b) DAISI (tuned)
(c) DAISI ($\epsilon=0$)
Figure 3: A comparison of filtering results from bootstrap particle filter (BPF) vs DAISI on the L63 system. We display the ground truth alongside the filtering mean and 99% credible interval.
(a) Truth
(b) Noisy
(c) Sparse
(d) Multimodal
(e) Saturating
Figure 4: Ensemble mean and a single member for DAISI and LETKF for the SQG experiments at the last step of the assimilated trajectory.
We evaluate DAISI on the Lorenz ’63 (L63) system to assess its filtering performance. As discussed in Section 3.1, DAISI does not sample exactly from the filtering distribution; our goal here is to demonstrate how an appropriate choice of $t_{\min}$ and $\epsilon$ can substantially reduce errors. The L63 system is integrated with standard parameters for 5000 RK4 time steps from random initial conditions, and the final 500 steps are used to test the assimilation methods. For reference, we compare DAISI with the Bootstrap Particle Filter (BPF) (gordon1993novel), a gold standard for low-dimensional filtering. As BPF is asymptotically exact, we take this as the “ground truth” filtering result to compare against. We also use an asymptotically exact guidance method for DAISI (see Appendix A.5.3) to isolate DAISI’s intrinsic error. For further details on the experimental setup, see Appendix D.3.
Table 1 summarizes the RMSE and probabilistic metrics—the Continuous Ranked Probability Score (CRPS) and Spread–Skill Ratio (SSR). For comparison, we also include the results of DAISI without inverse-sampling; this performs substantially worse than BPF across all metrics, demonstrating that incorporating dynamical information is essential for accurate filtering. With the default hyperparameters $t_{\min}\approx 0$ and $\epsilon=0$, DAISI still underperforms, and yields worse results than the version without inversion. Tuning $t_{\min}$ alone [^1] The hyperparameters are tuned on a short held-out validation trajectory to minimize CRPS. already leads to substantial improvement, and jointly tuning $t_{\min}$ and $\epsilon$ produces the best performance, achieving metrics closest to those of BPF.
Visually, we see that after tuning both $t_{\min}$ and $\epsilon$, the DAISI filtering results (Figure 3) closely resemble those from BPF (Figure 3). To see the effect of $\epsilon$, we repeat the experiment with the same tuned $t_{\min}$ but fix $\epsilon=0$. The results, shown in Figure 3 and summarized in the final row of Table 1, exhibit similar mean behaviour but has narrower uncertainty bands. This reduction in ensemble spread leads to a much lower SSR and a degraded CRPS, despite similar RMSE. This highlights the importance of $\epsilon$ for maintaining ensemble spread, while tuning $t_{\min}$ is necessary for accuracy.
4.2 Surface Quasi-Geostrophic (SQG) Dynamics
We next evaluate DAISI on a two-layer Surface Quasi-Geostrophic (SQG) model, a standard benchmark for turbulent geophysical flows. The system evolves potential temperature on two horizontal surfaces confining a fluid. Despite its simplicity, SQG exhibits strong sensitivity to initial conditions and exhibit multiscale turbulent behavior while remaining computationally tractable, making it a valuable benchmark for DA (tulloch2009quasigeostrophic).
Following (liang2025ensemble), we use a $64\times 64$ grid over 100 time steps with 3-hour intervals, and additionaly include a single $256\times 256$ experiment to assess scalability. DAISI is tested across a range of observation operators, noise levels, and sparsity settings listed in Table 2. Since the guidance methods in Section 2.3 have been applied almost exclusively to Gaussian observations, we stay within this setting, but note that this is not a limitation with DAISI itself.
We compare DAISI to both classical and ML-based DA methods. Classical baselines include the Localized Ensemble Transport Filter (LETKF), while ML baselines include FlowDAS and the Ensemble Score Filter (EnSF). DAISI uses a U-Net architecture based on Karras2022edm with 3.5M parameters. With the exception of FlowDAS, we initialize all methods from ${\bm{x}}_{0}\sim\mathcal{N}({\bm{x}}_{0}^{\mathrm{gt}},\sigma_{\mathrm{init}}^{2}\mathbf{I})$, $\sigma_{\mathrm{init}}=3$, and propagate ensemble members using the numerical model. FlowDAS instead uses its own learned autoregressive forecast model conditioned on the six previous states. Therefore assimilation begins from samples drawn from $p({\bm{x}}_{0}\mid{\bm{x}}_{-1}^{\mathrm{gt}},\dots,{\bm{x}}_{-6}^{\mathrm{gt}})$. All methods use 20 members.
Table 2: Experiment configurations.
Experiment Data Obs operator $\sigma_{\text{obs}}$ Sparsity
Noisy SQG-$64$ $x$ 55 $25\%$
Sparse SQG-$64$ $x$ 11 $5\%$
Multimodal SQG-$64$ $(x/7)^{2}$ 11 $25\%$
Saturating SQG-$64$ $\arctan(x)$ $0.01$ $25\%$
High-dim. SQG-$256$ $x$ 33 $5\%$
SEVIR SEVIR $x$ $0.001$ $10\%$
Table 3: The CRPS for experiments on SQG and SEVIR. We display the mean and standard deviation across 10 independent trajectories, averaged over the last 20 (10 for SEVIR) steps. The best score for each experiment is highlighted in bold and the second best with an underline.
Experim. DAISI LETKF FlowDAS EnSF
Noisy $\mathbf{1.32}{\scriptstyle\pm 0.13}$ $\underline{1.34}{\scriptstyle\pm 0.14}$ $2.81{\scriptstyle\pm 0.43}$ $5.15{\scriptstyle\pm 0.78}$
Sparse $\mathbf{1.73}{\scriptstyle\pm 0.18}$ $\underline{2.35}{\scriptstyle\pm 0.28}$ $3.34{\scriptstyle\pm 0.53}$ $4.20{\scriptstyle\pm 0.56}$
Multimo. $\mathbf{1.81}{\scriptstyle\pm 0.44}$ $\underline{1.97}{\scriptstyle\pm 0.69}$ $3.66{\scriptstyle\pm 0.42}$ $6.38{\scriptstyle\pm 0.75}$
Saturat. $\underline{1.54}{\scriptstyle\pm 0.09}$ $5.24{\scriptstyle\pm 0.35}$ $2.41{\scriptstyle\pm 0.39}$ $\mathbf{1.33}{\scriptstyle\pm 0.16}$
SEVIR $\mathbf{0.016}{\scriptstyle\pm 0.01}$ - $\underline{0.045}{\scriptstyle\pm 0.01}$ -
Table 3 summarizes CRPS across all SQG experiments. DAISI consistently achieves accurate assimilation, producing temporally coherent and physically plausible ensemble members (Figures 4 and 5). It matches LETKF in the noisy and high-dimensional settings, and clearly outperforms it under sparse or nonlinear observations. DAISI also remains stable when assimilation is performed every 12 hours instead of every 3, whereas LETKF degrades (Figure 12). In multimodal settings, DAISI reliably tracks all plausible modes, whereas LETKF collapses to a single mode and typically diverges once that mode becomes inconsistent (Figure 21).
For the saturating arctan observations, DAISI performs comparably to EnSF, which handles such nonlinearities well. However, EnSF fails in other regimes due to the lack of a learned prior, and is prone to mode collapse, requiring inflation to perform well. LETKF diverges under these nonlinear observations, and FlowDAS underperforms in all settings, reflecting its sensitivity to observations outside the support of its forecast model. Replacing the numerical model with the FlowDAS forecast model within DAISI shows no performance degradation, confirming FlowDAS’s failures arise from the filtering scheme rather than its forecast model (Table 11). Additional results are provided in Appendix D.4.5.
(a) High-dimensional
(b) SEVIR
Figure 5: Ensemble mean and a single member for DAISI and LETKF/FlowDAS for the High-dimensional/SEVIR experiments at the last step of the assimilated trajectory.
4.3 Precipitation Nowcasting using SEVIR
To evaluate on a real-world dataset, we apply DAISI to the Storm EVent Imagery and Radar (SEVIR) dataset (veillette2020sevir), a radar observation dataset of convective storms over the United States. We use the vertically integrated liquid on a $384\times$$384\text{\,}\mathrm{km}$ grid at $2\text{\,}\mathrm{km}$ resolution available every $10\text{\,}\mathrm{min}$ for $250\text{\,}\mathrm{min}$ (gao2023prediff).
This dataset has been applied to DA using the FlowDAS method in (chen2025flowdas). Thus, we mirror their setting and consider a linear Gaussian observation with standard deviation 0.001 and 10% sparsity. To generate the forecasts, we use the pre-trained forecasting model of (chen2025flowdas). This takes six consecutive frames as input and predicts the state 10 min into the future.
As shown in Figure 5, both DAISI and FlowDAS are able to accurately reconstruct the state, although the extreme events are better represented in DAISI. This is also reflected by the much lower CRPS, as shown in Table 3.
5 Conclusion
We introduced DAISI, a robust and flexible filtering framework built on flow-based generative priors. DAISI avoids retraining at every assimilation step by leveraging a fixed, pre-trained flow model and enables zero-shot conditional sampling for arbitrary observation operators. A key component is the inversion of the generative SDE: running the flow backward from the forecast ensemble recovers latent states that serve as informative initialization for conditional sampling, effectively combining the prior, the forecast model, and observational information in a unified way.
Empirically, DAISI delivers accurate filtering results across a spectrum of challenging settings, including sparse, noisy, nonlinear, and multimodal observations. It maintains physical coherence over long horizons, remains stable under infrequent assimilation, and outperforms classical approaches where they break down. Its modular design further supports seamless integration with both numerical simulators and modern ML-based forecasting models.
Overall, DAISI offers a flexible, robust, and expressive probabilistic data assimilation framework, capable of capturing complex posterior structures in high dimensions that are beyond the reach of classical filters.
6 Limitations & Future Work
A fundamental limitation of DAISI is that it does not sample exactly from the true filtering distribution. While our experiments show that tuning $t_{\min}$ and $\epsilon$ can mitigate this discrepancy, developing principled correction schemes for debiasing remains an important direction for future work.
DAISI also inherits the limitations of the underlying generative model. For the $64\times 64$ SQG experiment, our trained generative prior underestimated power across scales, causing degraded performance for the DAISI variant without inversion. Notably, DAISI itself appears robust to this issue, suggesting that the inversion step helps counteract spectral biases of the generative model; understanding this corrective effect more formally is an interesting open question.
Another limitation concerns the guidance mechanism. In this work, we use gradient-based guidance, which is known to introduce bias. Certain iterative schemes that repeatedly inject noise into the dynamics (e.g., (zhang2025improving)) are unlikely to work well in combination with DAISI.
From a computational standpoint, DAISI inherits the high sampling cost of flow-based generative models, driven by the high number of function evaluations required to integrate generative SDEs. Performing DA in latent spaces (andry2025appa) may offer avenues for reducing these costs.
Finally, the role of $\epsilon$ is richer than the limited setting considered here. More expressive parameterizations—such as matrix-valued or ensemble-adaptive noise levels—may offer additional flexibility and improved performance. We leave these extensions for future investigation.
Appendix A Details on Stochastic Interpolants
A.1 Stochastic Interpolants
Given a measure $\rho_{1}$, a (linear one-sided) stochastic interpolant is a stochastic process of the following form:
${\bm{z}}_{t}=\alpha_{t}{\bm{z}}_{1}+\beta_{t}{\bm{z}}_{0}$ (21)
where ${\bm{z}}_{1}\sim\rho_{1}$ and ${\bm{z}}_{0}\sim\mathcal{N}(\bm{0},\mathbf{I})$. We have the following result, found in Theorem 2.6 of (albergo2023stochastic).
Proposition A.1.
The probability distribution $\rho_{t}$ of the interpolant ${\bm{z}}_{t}$ admits Lebesgue densities $p(t)$ for all times $t\in[0,1]$ and moreover, it satisfies the endpoint conditions $\rho(0)=\mathcal{N}(\bm{0},\mathbf{I}),\rho(1)=\rho_{1}$. In addition, the Lebesgue densities satsify the transport equation
$\partial_{t}p_{t}+\nabla\cdot({\bm{b}}_{t}p_{t})=0,$ (22)
where ${\bm{b}}_{t}$ is the drift of the interpolant, defined by
${\bm{b}}(t,{\bm{z}}_{t})=\mathbb{E}[\dot{{\bm{z}}}_{t}|{\bm{z}}_{t}].$ (23)
This result implies that the flow map $\{\Phi_{t}\}_{t\in[0,1]}$ of the ODE
$\frac{\mathrm{d}{\bm{z}}_{t}}{\mathrm{d}t}={\bm{b}}(t,{\bm{z}}_{t}).$ (24)
transports $\mathcal{N}(\bm{0},\mathbf{I})$ to $\rho_{1}$, i.e., $(\Phi_{1})_{\sharp}\mathcal{N}(\bm{0},\mathbf{I})=\rho_{1}$.
Furthermore, (albergo2023stochastic) shows that the drift ${\bm{b}}(t,{\bm{z}})$ can be learned by minimizing the objective
$\mathcal{L}(\theta)$ $=\mathbb{E}_{{\bm{z}}_{0}\sim\mathcal{N}(\bm{0},\mathbf{I}),{\bm{z}}_{1}\sim\rho_{1},t\sim\mathcal{U}([0,1])}\Big[\|{\bm{{\bm{b}}}}_{\theta}(t,{\bm{z}}_{t})-(\dot{\alpha}_{t}{\bm{z}}_{1}+\dot{\beta}_{t}{\bm{z}}_{0})\|^{2}\Big].$ (25)
A.2 Turning ODEs into SDEs
We now show how one can transform an ODE into an SDE that shares the same marginal laws.
By Proposition A.1 we know that the marginals of the stochastic interpolant in Equation 21 satisfy the continuity equation
$\partial_{t}\rho_{t}+\nabla\cdot({\bm{b}}_{t}\rho_{t})=0.$ (26)
By noticing that for any non-negative $\epsilon_{t}$, we have the identity
$\epsilon_{t}\Delta p_{t}=\epsilon_{t}\nabla\cdot(p_{t}\nabla\log p_{t})=\nabla\cdot(\epsilon_{t}{\bm{s}}p_{t}),$ (27)
we can add and subtract a term from Equation 26, giving us the equivalent expression
$\partial_{t}p_{t}=-\nabla\cdot(({\bm{b}}+\epsilon_{t}{\bm{s}})p_{t})+\epsilon_{t}\Delta p_{t}.$ (28)
We recognize that this is the Fokker-Planck equation whose samples satisfies the forward/backward SDEs
$\mathrm{d}{\bm{z}}_{t}=({\bm{b}}(t,{\bm{z}}_{t})\pm\epsilon_{t}{\bm{s}}(t,{\bm{z}}_{t}))\mathrm{d}t+\sqrt{2\epsilon_{t}}\mathrm{d}W_{\pm t},$ (29)
${\bm{z}}_{0}\sim\rho_{0},\quad{\bm{z}}_{1}\sim\rho_{1},\quad t\in[0,1].$ (30)
Thus, we have recovered another family of generative models based on SDEs that share the same marginals as the generative ODE Equation 24.
A.3 Identities for drifts and score
We now derive some important relationships between the drift ${\bm{b}}(t,{\bm{z}})$, score ${\bm{s}}(t,{\bm{z}}):=\nabla\log p_{t}({\bm{z}})$, and $\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]$.
Taking the time derivative of (21), we have
$\dot{{\bm{z}}}_{t}=\dot{\alpha}_{t}{\bm{z}}_{1}+\dot{\beta}_{t}{\bm{z}}_{0}$ (31)
and therefore by definition of the drift, we have
${\bm{b}}(t,{\bm{z}})=\dot{\alpha}_{t}\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]+\dot{\beta}_{t}\mathbb{E}[{\bm{z}}_{0}|{\bm{z}}_{t}].$ (32)
Now, since ${\bm{z}}_{0}\sim\mathcal{N}(\bm{0},\mathbf{I})$, the interpolant expression implies that ${\bm{z}}_{t}|{\bm{z}}_{1}\sim\mathcal{N}(\alpha_{t}{\bm{z}}_{1},\beta_{t}^{2}I)$. Thus, we can write down the following Tweedie-type estimate for the score function
${\bm{s}}(t,{\bm{z}}_{t})$ $:=\nabla_{{\bm{z}}_{t}}\log p({\bm{z}}_{t})$ (33)
$=\frac{\nabla_{{\bm{z}}_{t}}p({\bm{z}}_{t})}{p({\bm{z}}_{t})}$ (34)
$=\frac{\nabla_{{\bm{z}}_{t}}\int p({\bm{z}}_{t}|{\bm{z}}_{1})p({\bm{z}}_{1})\mathrm{d}{\bm{z}}_{1}}{p({\bm{z}}_{t})}$ (35)
$=\frac{\int\nabla_{{\bm{z}}_{t}}p({\bm{z}}_{t}|{\bm{z}}_{1})p({\bm{z}}_{1})\mathrm{d}{\bm{z}}_{1}}{p({\bm{z}}_{t})}$ (36)
$=\frac{\int(\nabla_{{\bm{z}}_{t}}\log p({\bm{z}}_{t}|{\bm{z}}_{1}))p({\bm{z}}_{t}|{\bm{z}}_{1})p({\bm{z}}_{1})\mathrm{d}{\bm{z}}_{1}}{p({\bm{z}}_{t})}$ (37)
$=\int(\nabla_{{\bm{z}}_{t}}\log p({\bm{z}}_{t}|{\bm{z}}_{1}))p({\bm{z}}_{t}|{\bm{z}}_{1})\mathrm{d}{\bm{z}}_{1}.$ (38)
Using that ${\bm{z}}_{t}|{\bm{z}}_{1}\sim\mathcal{N}(\alpha_{t}{\bm{z}}_{1},\beta_{t}^{2}I)$, we get
$\nabla_{{\bm{z}}_{t}}\log p({\bm{z}}_{t}|{\bm{z}}_{1})=-\frac{{\bm{z}}_{t}-\alpha_{t}{\bm{z}}_{1}}{\beta_{t}^{2}},$ (39)
which allows us to obtain
${\bm{s}}(t,{\bm{z}}_{t})=-\frac{{\bm{z}}_{t}-\alpha_{t}\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]}{\beta_{t}^{2}}.$ (40)
Now, taking the conditional expectation $\mathbb{E}[\,\cdot\,|{\bm{z}}_{t}]$ of (21) we get that
${\bm{z}}_{t}=\alpha_{t}\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]+\beta_{t}\mathbb{E}[{\bm{z}}_{0}|{\bm{z}}_{t}],$ (41)
which implies
$\mathbb{E}[{\bm{z}}_{0}|{\bm{z}}_{t}]=-\beta_{t}{\bm{s}}(t,{\bm{z}}_{t}).$ (42)
Finally, combining these with the expression for $\bm{b}$} found earlier, we arrive at
${\bm{s}}(t,{\bm{z}}_{t})=\frac{\alpha_{t}{\bm{b}}(t,{\bm{z}}_{t})-\dot{\alpha}_{t}{\bm{z}}_{t}}{\beta_{t}\gamma_{t}},$ (43)
and
$\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]=\frac{\beta_{t}{\bm{b}}(t,{\bm{z}}_{t})-\dot{\beta}{\bm{z}}_{t}}{\gamma_{t}},$ (44)
where $\gamma_{t}:=\dot{\alpha}_{t}\beta_{t}-\alpha_{t}\dot{\beta}_{t}$.
A.4 Conditional drift and score
Now replace the data distribution $p_{1}$ with the posterior distribution
$p_{1}^{{\bm{y}}}({\bm{z}}_{1}):=\frac{p({\bm{y}}|{\bm{z}}_{1})p_{1}({\bm{z}}_{1})}{p({\bm{y}})},$ (45)
and again consider a stochastic interpolant (21), where now ${\bm{z}}_{1}\sim p_{1}^{{\bm{y}}}$. Then the law of the interpolant is given by
$p_{t}^{{\bm{y}}}({\bm{z}}_{t})$ $:=\int p({\bm{z}}_{t}|{\bm{z}}_{1})p_{1}^{{\bm{y}}}({\bm{z}}_{1})\mathrm{d}{\bm{z}}_{1}$ (46)
$=\frac{1}{p({\bm{y}})}\int p({\bm{z}}_{t}|{\bm{z}}_{1})p_{1}({\bm{z}}_{1})p({\bm{y}}|{\bm{z}}_{1})\mathrm{d}{\bm{z}}_{1}$ (47)
$=\frac{1}{p({\bm{y}})}\int p({\bm{z}}_{1}|{\bm{z}}_{t})p_{t}({\bm{z}}_{t})p({\bm{y}}|{\bm{z}}_{1})\mathrm{d}{\bm{z}}_{1}$ (48)
$=\frac{1}{p({\bm{y}})}p_{t}({\bm{z}}_{t})\int p({\bm{z}}_{1}|{\bm{z}}_{t})p({\bm{y}}|{\bm{z}}_{1})\mathrm{d}{\bm{z}}_{1}$ (49)
$=\frac{1}{p({\bm{y}})}p_{t}({\bm{z}}_{t})p({\bm{y}}|{\bm{z}}_{t}),$ (50)
where $p_{t}({\bm{z}}_{t})$ is the law of the unconditional interpolant. We note that this derivation relies on the fact that the interpolant ${\bm{z}}_{t}$ has the same conditional structure $p({\bm{z}}_{t}|{\bm{z}}_{s})$, regardless of whether the target is $p_{1}({\bm{z}}_{1})$ or $p_{1}^{\bm{y}}({\bm{z}}_{1})$.
Thus, the conditional score is given by
${\bm{s}}^{{\bm{y}}}(t,{\bm{z}}_{t})$ $:=\nabla_{{\bm{z}}_{t}}\log p_{t}^{{\bm{y}}}({\bm{z}}_{t})$ (51)
$=\nabla_{{\bm{z}}_{t}}\log p_{t}({\bm{z}}_{t})+\nabla_{{\bm{z}}_{t}}\log p({\bm{y}}|{\bm{z}}_{t})$ (52)
$={\bm{s}}(t,{\bm{z}}_{t})+\nabla_{{\bm{z}}_{t}}\log p({\bm{y}}|{\bm{z}}_{t}),$ (53)
where ${\bm{s}}(t,{\bm{z}}_{t})$ denotes the score of the interpolant that samples from the original measure $p_{1}$. Next, using our relation between score and drift, which holds for arbitrary data measures and therefore also the case of sampling from the posterior, we have the following corresponding drift
${\bm{b}}^{{\bm{y}}}(t,{\bm{z}}_{t})$ $=\frac{\dot{\alpha}_{t}{\bm{z}}_{t}+\beta_{t}\gamma_{t}{\bm{s}}^{{\bm{y}}}(t,{\bm{z}}_{t})}{\alpha_{t}}$ (54)
$=\frac{\dot{\alpha}_{t}{\bm{z}}_{t}+\beta_{t}\gamma_{t}({\bm{s}}(t,{\bm{z}}_{t})+\nabla_{{\bm{z}}_{t}}\log p({\bm{y}}|{\bm{z}}_{t}))}{\alpha_{t}}$ (55)
$={\bm{b}}(t,{\bm{z}}_{t})+\lambda_{t}\nabla_{{\bm{z}}_{t}}\log p({\bm{y}}|{\bm{z}}_{t}),$ (56)
where $\lambda_{t}:=\beta_{t}\gamma_{t}/\alpha_{t}$, and ${\bm{b}}(t,{\bm{z}}_{t})$ denotes the drift corresponding to the original interpolant. Then, by Proposition A.1, the flow of the ODE with drift ${\bm{b}}^{{\bm{y}}}$ transports $\mathcal{N}(\bm{0},\mathbf{I})$ to $p_{1}^{{\bm{y}}}$. This observation serves as the basis for sampling from the posterior using guidance, which we look at in the following section.
A.5 Guidance Methods
As demonstrated in the previous section, conditional sampling requires access to the likelihood score $\nabla_{{\bm{z}}_{t}}\log p({\bm{y}}|{\bm{z}}_{t})$. Unfortunately, this is, for the most part, analytically intractable and requires approximations. In this section, we present three such approximations that are used in this work. In addition to the approximation below, we may also multiply the likelihood score with a guidance strength $\zeta>0$, which we tune based on the problem setup.
A.5.1 Diffusion Posterior Sampling (DPS)
DPS (chung2023diffusion) proceeds by approximating $p({\bm{y}}\mid{\bm{z}}_{t})=\mathbb{E}_{{\bm{z}}_{1}|{\bm{z}}_{t}}[p({\bm{y}}\mid{\bm{z}}_{1})]$ by simply taking the expectation inside the likelihood:
$p({\bm{y}}\mid{\bm{z}}_{t})\approx p({\bm{y}}\mid\hat{{\bm{z}}}_{1}),$ (57)
where $\hat{{\bm{z}}}_{1}({\bm{z}}_{t})=\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]$. This incurs a bias, known as the Jensen gap; however, in many inverse problem settings, the approximation is known to work well. The resulting likelihood score used in DPS is thus
$\nabla_{{\bm{z}}_{t}}\log p({\bm{y}}\mid{\bm{z}}_{t})\approx\nabla_{{\bm{z}}_{t}}\log p({\bm{y}}\mid\hat{{\bm{z}}}_{1}({\bm{z}}_{t})),$ (58)
where we estimate $\hat{{\bm{z}}}_{1}({\bm{z}}_{t})=\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]$ explicitly from the drift via the relation (44).
When $p({\bm{y}}\mid{{\bm{z}}}_{1})$ is a Gaussian, we use the additional scaling factor proposed by (chung2023diffusion) which gives the expression
$\nabla_{{\bm{z}}_{t}}\log p({\bm{y}}\mid{\bm{z}}_{t})\approx\nabla_{{\bm{z}}_{t}}^{\top}\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]\mathbf{H}_{t}^{\top}({\bm{y}}-\mathcal{H}(\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]))$ (59)
where $\mathbf{H}_{t}:=\nabla_{{\bm{z}}_{1}}\mathcal{H}({\bm{z}}_{1})|_{{\bm{z}}_{1}=\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]}$.
A.5.2 Moment-Matching Posterior Sampling (MMPS)
MMPS extends DPS by considering the approximation $p({\bm{z}}_{1}|{\bm{z}}_{t})\approx\mathcal{N}(\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}],\mathbb{V}[{\bm{z}}_{1}|{\bm{z}}_{t}])$, where $\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]$ is obtained as before, and $\mathbb{V}[{\bm{z}}_{1}|{\bm{z}}_{t}]:=\mathbb{E}[{\bm{z}}_{1}{\bm{z}}_{1}^{\top}|{\bm{z}}_{t}]-\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]^{\top}$ can be computed using the following formula (Rozet_Andry_Lanusse_Louppe_2024):
$\mathbb{V}[{\bm{z}}_{1}|{\bm{z}}_{t}]$ $=\frac{\beta_{t}^{2}}{\alpha_{t}}\nabla_{{\bm{z}}_{t}}^{\top}\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}].$ (60)
Subsequently, the likelihood score can be approximated as
$\nabla_{{\bm{z}}_{t}}\log p({\bm{y}}|{\bm{z}}_{t})$ $=\nabla_{{\bm{z}}_{t}}\log\left(\int_{\mathbb{R}^{d}}p({\bm{y}}|{\bm{z}}_{1})p({\bm{z}}_{1}|{\bm{z}}_{t})\mathrm{d}{\bm{z}}_{1}\right)$ (61)
$\approx\nabla_{{\bm{z}}_{t}}\log\left(\int_{\mathbb{R}^{d}}p({\bm{y}}|{\bm{z}}_{1})\mathcal{N}({\bm{z}}_{1}\mid\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}],\,\mathbb{V}[{\bm{z}}_{1}|{\bm{z}}_{t}])\mathrm{d}{\bm{z}}_{1}\right)$ (62)
$\approx\nabla_{{\bm{z}}_{t}}^{\top}\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]\mathbf{H}_{t}^{\top}\left(\sigma_{{\bm{y}}}^{2}\mathbf{I}+\frac{\beta_{t}^{2}}{\alpha_{t}}\mathbf{H}_{t}\nabla_{{\bm{z}}_{t}}^{\top}\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}])\mathbf{H}_{t}^{\top}\right)^{-1}\big({\bm{y}}-\mathcal{H}(\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}])\big),$ (63)
where $\mathbf{H}_{t}:=\nabla_{{\bm{z}}_{1}}\mathcal{H}({\bm{z}}_{1})|_{{\bm{z}}_{1}=\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]}$. The last line becomes exact when $p({\bm{y}}|{\bm{z}}_{1})=\mathcal{N}({\bm{y}}\mid\mathcal{H}{\bm{z}}_{1},\sigma_{{\bm{y}}}^{2}\mathbf{I})$, where $\mathcal{H}$ is a linear operator.
A.5.3 Monte Carlo Guidance
On problems with smaller dimensions, we can develop an asymptotically exact guidance method using Monte Carlo integration. This follows from the following straightforward computation
$\nabla_{{\bm{z}}_{t}}\log p({\bm{y}}|{\bm{z}}_{t})$ $=\nabla_{{\bm{z}}_{t}}\log\left(\int_{\mathbb{R}^{d}}p({\bm{y}}|{\bm{z}}_{1})p({\bm{z}}_{1}|{\bm{z}}_{t})\mathrm{d}{\bm{z}}_{1}\right)$ (64)
$=\nabla_{{\bm{z}}_{t}}\log\left(\int_{\mathbb{R}^{d}}p({\bm{y}}|{\bm{z}}_{1})\frac{p({\bm{z}}_{t}|{\bm{z}}_{1})p_{1}({\bm{z}}_{1})}{\int_{\mathbb{R}^{d}}p({\bm{z}}_{t}|{\bm{z}}_{1})p_{1}({\bm{z}}_{1})\mathrm{d}{\bm{z}}_{1}}\mathrm{d}{\bm{z}}_{1}\right)$ (65)
$\stackrel{{\scriptstyle\text{MC}}}{{\approx}}\nabla_{{\bm{z}}_{t}}\log\left(\frac{\sum_{i=1}^{J}p({\bm{y}}|{\bm{z}}_{1}^{(i)})p({\bm{z}}_{t}|{\bm{z}}_{1}^{(i)})}{\sum_{j=1}^{J}p({\bm{z}}_{t}|{\bm{z}}_{1}^{(j)})}\right),\quad\text{for}\quad{\bm{z}}_{1}^{(i)},{\bm{z}}_{1}^{(j)}\sim p_{1}.$ (66)
Using the formulation of the stochastic interpolant, we have $p({\bm{z}}_{t}|{\bm{z}}_{1})=\mathcal{N}({\bm{z}}_{t}|\alpha_{t}{\bm{z}}_{1},\beta_{t}^{2}I)$ and provided we have a closed form expression for $p({\bm{y}}|{\bm{z}}_{1})$, we can compute the log-density $\log p({\bm{y}}|{\bm{z}}_{t})$ via Monte Carlo and use automatic differentiation to calculate its ${\bm{z}}_{t}$-gradient.
A.6 Rescaling of trained interpolant
Often, the model is trained using normalized data ${\bm{w}}_{t}=({\bm{z}}_{t}-\mu)/\sigma$ to stabilize training. If we have an interpolant in the normalized space,
${\bm{w}}_{t}=\alpha_{t}{\bm{w}}_{1}+\beta_{t}\bm{\xi},\quad\bm{\xi}\sim\mathcal{N}(\bm{0},\mathbf{I}),$ (67)
then, in the original space, we have the following un-normalized interpolant and its time derivative:
${\bm{z}}_{t}$ $=\varphi({\bm{w}}_{t}):=\sigma{\bm{w}}_{t}+\mu$ (68)
$=\alpha_{t}{\bm{z}}_{1}+\sigma\beta_{t}\bm{\xi}+(1-\alpha_{t})\mu,$ (69)
$\dot{{\bm{z}}}_{t}$ $=\dot{\alpha}_{t}{\bm{z}}_{1}+\sigma\dot{\beta}_{t}\bm{\xi}-\dot{\alpha}_{t}\mu,$ (70)
where ${\bm{z}}_{1}=\sigma{\bm{w}}_{1}+\mu$. Noting that
[…]
$=\frac{1}{\sigma^{d}}\frac{1}{(2\pi\beta_{t}^{2})^{d/2}}\exp\left(-\frac{1}{2\beta_{t}^{2}}\left\|\left(\frac{{\bm{z}}_{t}-\mu}{\sigma}\right)-\alpha_{t}\left(\frac{{\bm{z}}_{1}-\mu}{\sigma}\right)\right\|^{2}\right)$ (72)
$=\frac{1}{\sigma^{d}}\frac{1}{(2\pi\beta_{t}^{2})^{d/2}}\exp\left(-\frac{1}{2\beta_{t}^{2}}\|{\bm{w}}_{t}-\alpha_{t}{\bm{w}}_{1}\|^{2}\right)$ (73)
[…]
and $\frac{\mathrm{d}\mathbb{P}_{Z}}{\mathrm{d}\mathbb{P}_{W}}=\mathrm{det}(D\varphi)=\sigma^{d}$, then by change of variables, we get
${\bm{b}}_{Z}(t,{\bm{z}}_{t})$ $=\int\dot{{\bm{z}}}_{1}\mathbb{P}(\mathrm{d}{\bm{z}}_{1}|{\bm{z}}_{t})$ (75)
$=\frac{\int\dot{{\bm{z}}}_{1}p({\bm{z}}_{t}|{\bm{z}}_{1})\mathbb{P}_{Z}(\mathrm{d}{\bm{z}}_{1})}{\int p({\bm{z}}_{t}|{\bm{z}}_{1})\mathbb{P}_{Z}(\mathrm{d}{\bm{z}}_{1})}$ (76)
$=\frac{\cancel{(\sigma^{d}/\sigma^{d})}\int(\sigma\dot{{\bm{w}}}_{1})p({\bm{w}}_{t}|{\bm{w}}_{1})\mathbb{P}_{W}(\mathrm{d}{\bm{w}}_{1})}{\cancel{(\sigma^{d}/\sigma^{d})}\int p({\bm{w}}_{t}|{\bm{w}}_{1})\mathbb{P}_{W}(\mathrm{d}{\bm{w}}_{1})}$ (77)
$=\sigma\mathbb{E}[\dot{{\bm{w}}_{1}}|{\bm{w}}_{t}]$ (78)
$=\sigma{\bm{b}}_{W}(t,({\bm{z}}_{t}-\mu)/\sigma).$ (79)
Similarly, for the score, we get
${\bm{s}}_{Z}(t,{\bm{z}}_{t})$ $=\nabla_{{\bm{z}}_{t}}\log p_{Z}({\bm{z}}_{t})$ (80)
$=D\varphi^{-1}({\bm{w}}_{t})\nabla_{{\bm{w}}_{t}}\Big(\log\underbrace{\frac{\mathrm{d}\mathbb{P}_{Z}}{\mathrm{d}\mathbb{P}_{W}}({\bm{w}}_{t})}_{=\sigma^{d}}+\log p_{W}({\bm{w}}_{t})\Big)$ (81)
$=\sigma^{-1}\nabla_{{\bm{w}}_{t}}\log p_{W}({\bm{w}}_{t})$ (82)
$=\sigma^{-1}{\bm{s}}_{W}(t,({\bm{z}}_{t}-\mu)/\sigma).$ (83)
From the relation (91), we can write down the scaled score in terms of the scaled drift as
${\bm{s}}_{Z}(t,{\bm{z}}_{t})=\frac{\alpha_{t}{\bm{b}}_{Z}(t,{\bm{z}}_{t})-\dot{\alpha}_{t}({\bm{z}}_{t}-\mu)}{\sigma^{2}\beta_{t}\gamma_{t}}.$ (84)
Similarly, we can express $\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]$ and $\mathbb{E}[\bm{\xi}|{\bm{z}}_{t}]$ in terms of ${\bm{b}}_{W}$:
$\mathbb{E}[{\bm{z}}_{1}\mid{\bm{z}}_{t}]$ $=\mu+\sigma\frac{\beta_{t}{\bm{b}}_{W}(t,\frac{{\bm{z}}_{t}-\mu}{\sigma})-\dot{\beta}_{t}\frac{{\bm{z}}_{t}-\mu}{\sigma}}{\gamma_{t}},$ (85)
$\mathbb{E}[\bm{\xi}\mid{\bm{z}}_{t}]$ $=-\frac{\alpha_{t}{\bm{b}}_{W}(t,\frac{{\bm{z}}_{t}-\mu}{\sigma})-\dot{\alpha}_{t}\frac{{\bm{z}}_{t}-\mu}{\sigma}}{\gamma_{t}}$ (86)
A.7 Probability flow ODE
Diffusion models work with SDEs instead of ODEs but can still be used within our framework through the so-called Probability Flow ODE (song2020score), which we demonstrate here.
Consider a forward (noising) SDE
$\mathrm{d}{\bm{z}}_{t}=f(t,{\bm{z}}_{t})\mathrm{d}t+g(t)\mathrm{d}W_{t}.$ (87)
This induces a reverse-time SDE given by
$\mathrm{d}{\bm{z}}_{t}=(f(t,{\bm{z}}_{t})-g(t)^{2}{\bm{s}}(t,{\bm{z}}_{t}))\mathrm{d}t+g(t)\mathrm{d}W_{t}$ (88)
which, after learning $\bm{s}$}, can be used to generate samples. The probability flow ODE sharing the same marginals as this SDE is given by
$\frac{\mathrm{d}{\bm{z}}_{t}}{\mathrm{d}t}=f(t,{\bm{z}}_{t})-\frac{1}{2}g(t)^{2}{\bm{s}}(t,{\bm{z}}_{t}).$ (89)
With the same argument as for the stochastic interpolant, this shares the same marginals as
$\mathrm{d}{\bm{z}}_{t}=\left(f(t,{\bm{z}}_{t})+\left(\epsilon_{t}-\frac{1}{2}g(t)^{2}\right){\bm{s}}(t,{\bm{z}}_{t})\right)\mathrm{d}t+\sqrt{2\epsilon_{t}}\mathrm{d}W_{t}.$ (90)
Appendix B Model Details
B.1 Design choices
In this work, we make the standard choice of a linear scheduler $\alpha_{t}=t,\beta_{t}=1-t$, which implies a cross-term $\gamma_{t}=1$ and the guided drift scaling $\lambda_{t}=(1-t)/t$. This gives the score
${\bm{s}}(t,{\bm{z}}_{t})=\frac{t{\bm{b}}(t,{\bm{z}}_{t})-{\bm{z}}_{t}}{1-t},$ (91)
and the expectation
$\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]={\bm{z}}_{t}+(1-t){\bm{b}}(t,{\bm{z}}_{t}).$ (92)
We note that this choice of schedule leads to an expectation given by a single Euler step to the final time. We also note that the score diverges as $t\to 1$. To avoid numerical issues, we let $\epsilon_{t}=\epsilon(1-t)$ for some $\epsilon\geq 0$. To solve the SDE, we use Euler–Maruyama; we do not use any additional Langevin correction steps, as these did not improve results.
Appendix C Entropy dissipation
In this section, we prove a Bakry-Émery-type entropy dissipation result (Proposition 3.1) (bakry2006diffusions), which describes the influence of the stochastic parameter $\epsilon$ on the information loss of the forecast ensembles when assimilating data with DAISI.
We restate the result below for convenience.
Proposition C.1.
Assume that the marginal laws $\{\rho_{t}\}_{t\in[0,1]}$ and $\{\rho_{t}^{\bm{y}}\}_{t\in[0,1]}$ of the interpolants bridging $\rho_{0}=\mathcal{N}(\bm{0},\mathbf{I})$ with $\rho_{1}=\mathbb{P}_{\infty}$ and $\rho_{1}=\mathbb{P}_{\infty}^{\bm{y}}$, respectively, satisfy a log-Sobolev inequality with constant $\lambda>0$, uniform in $t$. That is,
$\mathcal{KL}(\mathbb{P}||\rho_{t})\leq\frac{1}{2\lambda}\int_{\mathbb{R}^{d}}\left|\nabla\log\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\rho_{t}}({\bm{x}})\right|^{2}\mathbb{P}(\mathrm{d}{\bm{x}}),$ (93a)
$\mathcal{KL}(\mathbb{P}||\rho_{t}^{\bm{y}})\leq\frac{1}{2\lambda}\int_{\mathbb{R}^{d}}\left|\nabla\log\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\rho_{t}^{\bm{y}}}({\bm{x}})\right|^{2}\mathbb{P}(\mathrm{d}{\bm{x}}),$ (93b)
for any $t\in[0,1]$ and any probability measure $\mathbb{P}$. Then, we have
$\mathcal{KL}\left(\pi_{n,0,\epsilon}^{\text{DAISI}}\,||\,\mathbb{P}_{\infty}^{\bm{y}}\right)\leq e^{-4\lambda\epsilon}\,\mathcal{KL}\left(\hat{\pi}_{n}\,||\,\mathbb{P}_{\infty}\right),$ (94)
with equality at $\epsilon=0$.
Proof.
Consider a stochastic interpolant $\{{\bm{z}}_{t}\}$ bridging between $\rho_{0}=\mathcal{N}(\bm{0},\mathbf{I})$ and $\rho_{1}=\mathbb{P}_{\infty}$ and let $p_{t}$ denote the Lebesgue density of $\rho_{t}=\mathrm{Law}({\bm{z}}_{t})$. Then, we know from Proposition A.1 that $\{p_{t}\}_{t\in[0,1]}$ is a solution to the transport equation
$\frac{\partial p_{t}}{\partial t}({\bm{x}})+\mathrm{div}({\bm{b}}_{t}({\bm{x}})p_{t}({\bm{x}}))=0,\quad p_{0}({\bm{x}})=\frac{\mathrm{d}\rho_{0}}{\mathrm{d}{\bm{x}}},$ (95)
where we denoted by $\mathrm{d}{\bm{x}}$ the Lebesgue measure on $\mathbb{R}^{d}$. Now consider the backward SDE (up to time reparameterization)
$\mathrm{d}{\bm{z}}_{\tau}=-{\bm{b}}_{1-\tau}({\bm{x}})\mathrm{d}\tau+\epsilon\nabla\log p_{1-\tau}({\bm{z}}_{\tau})+\sqrt{2\epsilon}\,\mathrm{d}W_{\tau},\quad z_{0}\sim\hat{\pi}_{n},$ (96)
solved for $\tau=0\rightarrow 1$, which is used in the inverse sampling step of DAISI. For simplicity, we also denote $q_{\tau}:=p_{1-\tau}$, which satisfies
$\frac{\partial q_{\tau}}{\partial\tau}({\bm{x}})=\mathrm{div}({\bm{b}}_{1-\tau}({\bm{x}})q_{\tau}({\bm{x}})),\quad q_{0}({\bm{x}})=\frac{\mathrm{d}\mathbb{P}_{\infty}}{\mathrm{d}{\bm{x}}}({\bm{x}}),$ (97)
by (95). One can check that the Fokker-Planck equation of the SDE (96) reads
$\frac{\partial r_{\tau}}{\partial\tau}({\bm{x}})=\mathrm{div}\Big({\bm{b}}_{1-\tau}({\bm{x}})r_{\tau}({\bm{x}})+\epsilon\,r_{\tau}\nabla\log(r_{\tau}({\bm{x}})/q_{\tau}({\bm{x}}))\Big),\quad r_{0}({\bm{x}})=\frac{\mathrm{d}\hat{\pi}_{n}}{\mathrm{d}{\bm{x}}}({\bm{x}})$ (98)
and consider the evolution of the Kullback-Leibler divergence between the distributions given by (97) and (98)
$\mathcal{KL}(r_{\tau}||q_{\tau})=\int_{\mathbb{R}^{d_{x}}}r_{\tau}({\bm{x}})\log\frac{r_{\tau}({\bm{x}})}{q_{\tau}({\bm{x}})}\mathrm{d}{\bm{x}}.$ (99)
Taking the time derivative of (99) yields
$\frac{\mathrm{d}}{\mathrm{d}\tau}\mathcal{KL}(r_{\tau}||q_{\tau})$ $=\int_{\mathbb{R}^{d_{x}}}\frac{\partial r_{\tau}}{\partial\tau}({\bm{x}})\log\frac{r_{\tau}({\bm{x}})}{q_{\tau}({\bm{x}})}\mathrm{d}{\bm{x}}+\int_{\mathbb{R}^{d_{x}}}r_{\tau}({\bm{x}})\frac{\partial}{\partial\tau}\left(\log\frac{r_{\tau}({\bm{x}})}{q_{\tau}({\bm{x}})}\right)\mathrm{d}{\bm{x}}$ (100)
[…]
$\qquad+\int_{\mathbb{R}^{d_{x}}}r_{\tau}({\bm{x}})\left(\frac{1}{r_{\tau}({\bm{x}})}\frac{\partial r_{\tau}}{\partial\tau}({\bm{x}})-\frac{1}{q_{\tau}({\bm{x}})}\frac{\partial q_{\tau}}{\partial\tau}({\bm{x}})\right)\mathrm{d}{\bm{x}}$ (101)
$=-\int_{\mathbb{R}^{d_{x}}}{\bm{b}}_{1-\tau}({\bm{x}})r_{\tau}({\bm{x}})\nabla\log\frac{r_{\tau}({\bm{x}})}{q_{\tau}({\bm{x}})}\mathrm{d}{\bm{x}}-\epsilon\int_{\mathbb{R}^{d_{x}}}r_{\tau}({\bm{x}})\left|\nabla\log\frac{r_{\tau}({\bm{x}})}{q_{\tau}({\bm{x}})}\right|^{2}\mathrm{d}{\bm{x}}+\cancel{\frac{\mathrm{d}}{\mathrm{d}\tau}\int_{\mathbb{R}^{d_{x}}}r_{\tau}({\bm{x}})\mathrm{d}{\bm{x}}}$
$\qquad-\int_{\mathbb{R}^{d_{x}}}\frac{r_{\tau}({\bm{x}})}{q_{\tau}({\bm{x}})}\mathrm{div}({\bm{b}}_{1-\tau}({\bm{x}})q_{\tau}({\bm{x}}))\mathrm{d}{\bm{x}},$ (102)
where we used integration-by-parts and (97) to arrive at the last line. To further simplify this expression, we note that
$-\int_{\mathbb{R}^{d_{x}}}{\bm{b}}_{1-\tau}({\bm{x}})r_{\tau}({\bm{x}})\nabla\log\frac{r_{\tau}({\bm{x}})}{q_{\tau}({\bm{x}})}\mathrm{d}{\bm{x}}$ $=-\int_{\mathbb{R}^{d_{x}}}{\bm{b}}_{1-\tau}({\bm{x}})r_{\tau}({\bm{x}})\nabla\log r_{\tau}({\bm{x}})\mathrm{d}{\bm{x}}+\int_{\mathbb{R}^{d_{x}}}{\bm{b}}_{1-\tau}({\bm{x}})r_{\tau}({\bm{x}})\nabla\log q_{\tau}({\bm{x}})\mathrm{d}{\bm{x}}$ (103)
$=-\int_{\mathbb{R}^{d_{x}}}{\bm{b}}_{1-\tau}({\bm{x}})\nabla r_{\tau}({\bm{x}})\mathrm{d}{\bm{x}}+\int_{\mathbb{R}^{d_{x}}}{\bm{b}}_{1-\tau}({\bm{x}})r_{\tau}({\bm{x}})\nabla\log q_{\tau}({\bm{x}})\mathrm{d}{\bm{x}}$ (104)
$=\int_{\mathbb{R}^{d_{x}}}r_{\tau}({\bm{x}})\mathrm{div}({\bm{b}}_{1-\tau}({\bm{x}}))\mathrm{d}{\bm{x}}+\int_{\mathbb{R}^{d_{x}}}{\bm{b}}_{1-\tau}({\bm{x}})r_{\tau}({\bm{x}})\nabla\log q_{\tau}({\bm{x}})\mathrm{d}{\bm{x}},$ (105)
and also
$-\int_{\mathbb{R}^{d_{x}}}\frac{r_{\tau}({\bm{x}})}{q_{\tau}({\bm{x}})}\mathrm{div}({\bm{b}}_{1-\tau}({\bm{x}})q_{\tau}({\bm{x}}))\mathrm{d}{\bm{x}}$ $=-\int_{\mathbb{R}^{d_{x}}}r_{\tau}({\bm{x}})\mathrm{div}({\bm{b}}_{1-\tau}({\bm{x}}))\mathrm{d}{\bm{x}}-\int_{\mathbb{R}^{d_{x}}}\frac{r_{\tau}({\bm{x}}){\bm{b}}_{1-t}({\bm{x}})}{q_{\tau}({\bm{x}})}\nabla q_{\tau}({\bm{x}})\mathrm{d}{\bm{x}}$ (106)
$=-\int_{\mathbb{R}^{d_{x}}}r_{\tau}({\bm{x}})\mathrm{div}({\bm{b}}_{1-\tau}({\bm{x}}))\mathrm{d}{\bm{x}}-\int_{\mathbb{R}^{d_{x}}}{\bm{b}}_{1-\tau}({\bm{x}})r_{\tau}({\bm{x}})\nabla\log q_{\tau}({\bm{x}})\mathrm{d}{\bm{x}}.$ (107)
Hence, these two terms cancel out, giving us
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{KL}(r_{\tau}||q_{\tau})=-\epsilon\int_{\mathbb{R}^{d_{x}}}r_{\tau}({\bm{x}})\left|\nabla\log\frac{r_{\tau}({\bm{x}})}{q_{\tau}({\bm{x}})}\right|^{2}\mathrm{d}{\bm{x}}.$ (108)
Now, from our assumption that $\{p_{t}\}_{t\in[0,1]}$ and therefore $\{q_{\tau}\}_{\tau\in[0,1]}$ satisfies the log-Sobolev inequality (93a), we obtain the estimate
$\frac{\mathrm{d}}{\mathrm{d}\tau}\mathcal{KL}(r_{\tau}||q_{\tau})\leq-2\lambda\epsilon\mathcal{KL}(r_{\tau}||q_{\tau}).$ (109)
Then, applying Grönwall’s inequality, we arrive at
$\mathcal{KL}(r_{\tau}||q_{\tau})$ $\leq e^{-2\lambda\epsilon\tau}\mathcal{KL}(r_{0}||q_{0})$ (110)
$=e^{-2\lambda\epsilon\tau}\mathcal{KL}(\hat{\pi}_{n}||\mathbb{P}_{\infty}),$ (111)
where we slightly abused notation and used the same notation $\mathcal{KL}(\,\cdot\,||\,\cdot\,)$ to denote the Kullback-Leibler divergences between two measures and their corresponding Lebesgue densities. Now, taking the limit $\tau\rightarrow 1$, we get
[…]
where we note that $\rho_{0}(\mathrm{d}{\bm{x}})=q_{1}({\bm{x}})\mathrm{d}{\bm{x}}$ and we defined $\hat{\rho}_{0}^{\epsilon}(\mathrm{d}{\bm{x}}):=r_{1}({\bm{x}})\mathrm{d}{\bm{x}}$, which may be thought of as the latent representation of the predictive distribution $\hat{\pi}$ in noise space.
Next, we apply an identical argument for the guided sampling step of DAISI. That is, we now consider an interpolant $\{{\bm{z}}_{t}\}_{t\in[0,1]}$ bridging between $\rho_{0}=\mathcal{N}(\bm{0},\mathbf{I})$ and $\rho_{1}=\mathbb{P}_{\infty}^{\bm{y}}$. As per discussion in Section A.4, the corresponding law $p_{t}^{\bm{y}}$ satisfies the transport equation
$\frac{\partial p_{t}^{\bm{y}}}{\partial t}({\bm{x}})+\mathrm{div}({\bm{b}}_{t}^{\bm{y}}({\bm{x}})p_{t}^{\bm{y}}({\bm{x}}))=0,\quad p_{0}^{\bm{y}}({\bm{x}})=\frac{\mathrm{d}\rho_{0}}{\mathrm{d}{\bm{x}}}({\bm{x}}),$ (113)
with ${\bm{b}}_{t}^{\bm{y}}$ defined in (56). Now, consider the forward guided SDE
$\mathrm{d}{\bm{z}}_{t}={\bm{b}}^{\bm{y}}_{t}({\bm{x}})\mathrm{d}t+\epsilon\nabla\log p^{\bm{y}}_{t}({\bm{z}}_{t})+\sqrt{2\epsilon}\,\mathrm{d}W_{t},\quad z_{0}\sim\hat{\rho}_{0}^{\epsilon},$ (114)
whose marginal laws can be checked to solve the following Fokker-Planck equation
$\frac{\partial r_{t}^{\bm{y}}}{\partial t}({\bm{x}})=\mathrm{div}\Big(-{\bm{b}}_{t}^{\bm{y}}({\bm{x}})r_{t}^{\bm{y}}({\bm{x}})+\epsilon\,r_{t}^{\bm{y}}\nabla\log(r_{t}^{\bm{y}}({\bm{x}})/p_{t}^{\bm{y}}({\bm{x}}))\Big),\quad r_{0}^{\bm{y}}({\bm{x}})=\frac{\mathrm{d}\hat{\rho}_{0}^{\epsilon}}{\mathrm{d}{\bm{x}}}({\bm{x}}).$ (115)
Then, by a similar computation as before, we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{KL}(r_{t}^{\bm{y}}||p_{t}^{\bm{y}})=-\epsilon\int_{\mathbb{R}^{d_{x}}}r_{t}^{\bm{y}}({\bm{x}})\left|\nabla\log\frac{r_{t}^{\bm{y}}({\bm{x}})}{p_{t}^{\bm{y}}({\bm{x}})}\right|^{2}\mathrm{d}{\bm{x}},$ (116)
and again assuming that the conditional laws $\{p_{t}^{\bm{y}}\}_{t\in[0,1]}$ satisfy the uniform log-Sobolev inequality (93b) with uniform constant $\lambda$, we get
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{KL}(r_{t}^{\bm{y}}||p_{t}^{\bm{y}})=-2\lambda\epsilon\mathcal{KL}(r_{t}^{\bm{y}}||p_{t}^{\bm{y}})$ (117)
[…]
Taking the limit $t\rightarrow 1$, we get
$\mathcal{KL}(\pi_{n,0,\epsilon}^{\text{DAISI}}||\mathbb{P}_{\infty}^{\bm{y}})$ $\leq e^{-2\lambda\epsilon}\mathcal{KL}(\hat{\rho}_{0}^{\epsilon}||\rho_{0})$ (119)
[…]
which proves the inequality (112).
Now, from (108) and (116), we also see that when $\epsilon=0$, we have that
$\mathcal{KL}(\pi_{n,0,0}^{\text{DAISI}}||\mathbb{P}_{\infty}^{\bm{y}})=\mathcal{KL}(\tilde{\rho}_{0}^{0}||\rho_{0})=\mathcal{KL}(\hat{\pi}_{n}||\mathbb{P}_{\infty}),$ (121)
which completes the proof. ∎
Remark C.2.
Inequality (112) describes how DAISI exponentially contracts the KL divergence to the posterior measure $\mathbb{P}_{\infty}^{\bm{y}}$ as $\epsilon\rightarrow\infty$. In other words, this shows how noise in the generative SDE progressively “erases” the discrepancy between the forecast measure $\hat{\pi}_{n}$ and the background measure $\mathbb{P}_{\infty}$. On the other extreme, when $\epsilon=0$, this discrepancy (i.e., information about the forecast) is exactly preserved. We also note that the proof relies on the assumption that $t_{\min}=0$, however, we expect similar behaviour to hold for $t_{\min}>0$.
Remark C.3.
The uniform log-Sobolev assumption (93) is a strong regularity assumption, indicating that the interpolant path $\{\rho_{t}\}_{t\in[0,1]}$ is smooth and does not have any irregularities, e.g., develop narrow spikes or heavy tails. While this is normally an unreasonable assumption for high-dimensional distributions arising in generative modelling, the information loss behaviour as $\epsilon\rightarrow\infty$ can be observed empirically.
Appendix D Experimental details
D.1 Metrics
Given an assimilated state $\hat{X}$ at step $t$ we define the Root Mean Squared Error (RMSE) as
$\text{RMSE}^{t}=\sqrt{\frac{1}{|G||D|}\sum_{g\in G}\sum_{d\in D}(\bar{\hat{X}}_{g,d}^{t}-X_{g,d}^{t})^{2}},$ (122)
where $G$ is the set of grid points, $D$ is the set of variables at each grid point, and
$\bar{\hat{X}}^{t}=\frac{1}{N_{ens}}\sum^{N_{ens}}_{ens=1}\hat{X}^{t,ens}$ (123)
is the ensemble mean of the assimilated state. We also evaluate the RMSE of each ensemble member ensens
$\text{RMSE}^{t,ens}=\sqrt{\frac{1}{|G||D|}\sum_{g\in G}\sum_{d\in D}(\hat{X}_{g,d}^{t,ens}-X_{g,d}^{t})^{2}},$ (124)
where $\hat{X}^{t,ens}$ is the assimilated state at time $t$ for ensemble member ensens.
To measure the calibration of the ensemble forecasts, we measure the Continuous Ranked Probability Score (CRPS) (gneiting2007strictly). We follow and compute the fair unbiased CRPS estimate (fairCRPS; zamo2018estimation) for our ensemble, which reads
[…]
Additionally, we measure the Spread Skill Ratio (SSR) to evaluate the ensemble calibration, where a well calibrated ensemble should have $\text{SSR}\approx 1$ (fortin2014should). This is defined as
$\text{SSR}^{t}=\frac{1}{|D|}\sum_{d\in D}\sqrt{\frac{N_{\text{ens}}+1}{N_{\text{ens}}}}\frac{\text{Spread}^{t}_{d}}{\text{RMSE}^{t}_{d}},$ (126)
where
$\text{Spread}^{t}_{d}=\sqrt{\frac{1}{|G|N_{\text{ens}}}\sum_{g\in G}\sum^{N_{ens}}_{ens=1}(\bar{\hat{X}}^{t,ens}_{g,d}-\hat{X}^{t,ens}_{g,d})^{2}}.$ (127)
In addition to ensuring that the ensemble mean and individual members are accurate and that the ensemble is well calibrated, we also require each ensemble member to be physically realistic and to reproduce the energy spectrum of the ground truth. To evaluate this, we compute the Log Spectral Distance (LSD). Given a field $X$ defined on a 2D grid, we define its 2D power spectrum as
$P_{X}(\mathbf{k})=\left|\mathcal{F}\{X\}(\mathbf{k})\right|^{2},$ (128)
where $\mathcal{F}\{\cdot\}$ denotes the 2D Fourier transform and $\mathbf{k}=(k_{x},k_{y})$ is the wavenumber vector. We obtain an isotropic (radially averaged) power spectrum by averaging over all Fourier coefficients with the same radial wavenumber $r=\|\mathbf{k}\|$:
$\bar{P}_{X}(r)=\frac{1}{N(r)}\sum_{\|\mathbf{k}\|=r}P_{X}(\mathbf{k}),$ (129)
where $N(r)$ is the number of coefficients at radius $r$.
The Log Spectral Distance (LSD) between a forecast field $\hat{X}$ and the ground truth $X$ at time $t$ is then defined as
$\text{LSD}^{t}=\sqrt{\frac{1}{R}\sum_{r=1}^{R}\left(\log\!\left(\bar{P}_{\hat{X}}(r)+\varepsilon\right)-\log\!\left(\bar{P}_{X}(r)+\varepsilon\right)\right)^{2}},$ (130)
where $R$ is the maximum resolved radial wavenumber and $\varepsilon$ is a small constant included for numerical stability.
When we have access to samples $\{{\bm{x}}_{i}\}_{i=1}^{N}$, $\{{\bm{y}}_{j}\}_{j=1}^{M}$ from two measures $\pi_{1}$ and $\pi_{2}$, respectively, we can compute a particular notion of distance between them, using the Maximum Mean Discrepancy (MMD) (gretton2012kernel), defined as the distance between the kernel embeddings of the two measures in the corresponding reproducing kernel Hilbert space (RKHS) $\mathcal{H}$. In practice, given a kernel $k(\cdot,\cdot):\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, this is computed as
$\text{MMD}^{2}(\pi_{1},\pi_{2})$ $\approx\left\|\frac{1}{N}\sum_{i=1}^{N}k({\bm{x}}_{i},\cdot)-\frac{1}{M}\sum_{j=1}^{M}k({\bm{y}}_{j},\cdot)\right\|_{\mathcal{H}}^{2}$ (131)
$=\frac{1}{N(N-1)}\sum_{i=1}^{N}\sum_{j\neq i}^{N}k({\bm{x}}_{i},{\bm{x}}_{j})-\frac{2}{NM}\sum_{i=1}^{N}\sum_{j=1}^{M}k({\bm{x}}_{i},{\bm{y}}_{j})+\frac{1}{M(M-1)}\sum_{i=1}^{M}\sum_{j=1}^{M}k({\bm{y}}_{i},{\bm{y}}_{j}).$ (132)
For the choice of $k(\cdot,\cdot)$, we use the squared-exponential kernel $k({\bm{x}},{\bm{y}})=\exp\left(-\|{\bm{x}}-{\bm{y}}\|^{2}/2\sigma^{2}\right)$, where the bandwidth $\sigma^{2}$ is chosen to be the median heuristic $\sigma^{2}=\mathrm{Median}(\{\|{\bm{x}}_{i}-{\bm{y}}_{j}\|\}_{i,j})/2$.
D.2 1D Gaussian Mixture
In our toy 1D experiment, we considered an artificial setup whereby the invariant measure is given by a 1D Gaussian mixture $\mathbb{P}_{\infty}(\cdot)=\sum_{k=1}^{K}\phi_{k}\,\mathcal{N}(\cdot|\mu_{k},\sigma_{k}^{2})$ with $K=3$ and
$(\phi_{1},\phi_{2},\phi_{3})=(0.5,0.3,0.2),\qquad(\mu_{1},\mu_{2},\mu_{3})=(0.0,3.0,-2.0),\qquad(\sigma_{1},\sigma_{2},\sigma_{3})=(1.0,0.5,0.8).$ (133)
We trained a stochastic interpolant model on $50,000$ i.i.d. samples of this Gaussian mixture model, where we used a standard MLP with one hidden layer of size 64 and ReLU activations to parameterize the drift ${\bm{b}}({\bm{x}},t)$.
We use this toy experiment to understand the effect of the inherent error of DAIS on sampling from the filtering distribution, as highlighted in Section 3.1, and to investigate the impact of the hyperparameters $t_{\min}$ and $\epsilon$ on reducing this error.
To this end, we simulate the analysis step of DAISI by artificially constructing a predictive distribution $\hat{\pi}\propto f\,\mathbb{P}_{\infty}$, where we took $f(x)=\exp\left(-\frac{1}{2\sigma^{2}}|x-\mu|^{2}\right)$ with $\mu=0.5$ and $\sigma=1.5$. We obtained samples $\{\hat{x}_{i}\}_{i=1}^{N}$ from $\hat{\pi}$ via samples $\{x_{i}\}_{i=1}^{N}$ of the Gaussian mixture $\mathbb{P}_{\infty}$ by first re-weighting the particles by $f$, then re-sampling according to the weights, i.e.,
  1. 1.
    Compute weights $w_{i}=f(x_{i})$ for all $i=1,\ldots,N$,
  2. 2.
    Resample particles $\hat{x}_{i}=x_{j_{i}}$, where $j_{i}\sim\mathrm{Multinomial}(w_{1},...,w_{N})$ for $i=1,\ldots,N$.
Then, starting from the particles $\{\hat{x}_{i}\}_{i=1}^{N}$, we applied the analysis step of DAISI to obtain the particles $\{x^{y}_{i}\}_{i=1}^{N}$ post-assimilation, where, for the observation model, we took $p(y|x)=\mathcal{N}(y|x,\sigma_{\text{obs}}^{2})$ with $y=2.5$ and $\sigma_{\text{obs}}=1.0$. For the ensemble size, we set $N=10,000$ to accurately capture the resulting distributions. For the guidance method, we considered the asymptotically exact Monte Carlo strategy (see Section A.5.3) to isolate the sources of error as arising solely from the inexactness of DAISI (there will still be errors arising from numerical discretization and Monte Carlo estimation; however, these errors can be made arbitrarily small). For the numerical discretisation of the interpolant ODE and SDE, we used Euler-Maruyama integration with $200$ time steps and for the Monte-Carlo integration used in guidance, we used $J=10,000$ particles. As a point of reference to compare against, we also obtain samples directly from the filtering distribution $\pi\propto p(y|\cdot)\hat{\pi}$ by a similar reweight-resample strategy used before to obtain samples from $\hat{\pi}$.
We simulated the analysis step of DAISI with various combinations of hyperparameters $t_{\min}$ and $\epsilon$ and display the heatmap of the MMD between samples of the filtering distribution $\pi$ and samples of the distribution $\pi^{\text{DAISI}}$ obtained by DAISI in Figure 7. We see that the best result is achieved by the combination $t_{\min}=0.3$ and $\epsilon=0.1$, which yields an MMD of 0.004, indicating a very close match with the true filtering distribution. We also display ablations with respect to the individual hyperparameters in Figure 2 (for ablation with respect to $t_{\min}$, with $\epsilon$ fixed to 0) and Figure 6 (for ablation with respect to $\epsilon$, with $t_{\min}$ fixed to $0.01$). The figures demonstrate that the “default” setting of $\epsilon=0$ and $t_{\min}\approx 0$ yields a distribution $\pi^{\text{DAISI}}$ that is much more “peaked” and therefore mismatched from the filter distribution $\pi$. This error can be reduced by increasing $t_{\min}$, which pulls the profile of $\pi^{\text{DAISI}}$ towards the predictive distribution $\hat{\pi}$, while increasing $\epsilon$ pulls it towards the posterior $\mathbb{P}_{\infty}^{y}$. Since the filtering distribution $\pi\propto f\mathbb{P}_{\infty}^{y}=p(y|\cdot)\hat{\pi}$ has elements of both $\mathbb{P}_{\infty}^{y}$ and $\hat{\pi}$, we find that a slight nudge in these directions can help to match $\pi$ better.
(c) $\quad\epsilon=0.0$
(d) $\quad\epsilon=0.1$
(e) $\quad\epsilon=1.0$
Figure 6: Ablation with respect to the $\epsilon$ hyperparameter in the 1D Gaussian mixture experiment, fixing $t_{\min}=0.01$. The distribution obtained by DAISI when $\epsilon=0$ is highly peaked, making the resulting distribution overconfident. By increasing $\epsilon$, it loses information about the predictive ensemble and pulls the distribution towards $\mathbb{P}_{\infty}^{y}$, which can help to rejuvenate sample variance.
Figure 7: Heatmap of the maximum mean discrepancy (MMD) between the true filtering distribution $\pi$ and the DAISI analysis $\pi^{\text{DAISI}}$ using different combinations of $t_{\min}$ and $\epsilon$. The highlighted square corresponds to the $(t_{\min},\epsilon)$-combination that yielded the lowest MMD. The corresponding distribution obtained by DAISI is plotted above (orange) against the true filtering distribution (blue). We observe a close match between the two distributions under the optimal hyperparameters.
D.3 Lorenz ’63
The Lorenz ’63 (L63) model (lorenz1963deterministic) is given by the following system of ODEs
$\frac{\mathrm{d}x}{\mathrm{d}t}$ $=\sigma(y-z)$ (134)
$\frac{\mathrm{d}y}{\mathrm{d}t}$ $=x(\rho-z)-y$ (135)
$\frac{\mathrm{d}z}{\mathrm{d}t}$ $=xy-\beta z$ (136)
with the parameters set to $\sigma=10$, $\rho=28$ and $\beta=\frac{8}{3}$. In our experiments in Section 4.1, we integrate this system forward in time using a fourth order Runge-Kutta solver with time step $\Delta t=0.01$.
D.3.1 Model training
To train the stochastic interpolant model, we generate training data $\{x_{i}\}_{i=0}^{N-1}$ by integrating the L63 ODE (134)–(136) starting from the initial condition $x_{0}=(0,1,1.05)$ with $N=10^{6}$. We consider an 80-20 split for training and validation. Both datasets are normalized using the empirical mean and standard deviation of the training portion:
$w_{i}=\frac{x_{i}-\mu_{\text{train}}}{\sigma_{\text{train}}}.$ (137)
To parameterize the drift ${\bm{b}}(w,t)$, we use an MLP with two hidden layers; each with 128 neurons:
$(w,t)\in\mathbb{R}^{4}\;\mapsto\;128\;\mapsto\;128\;\mapsto\;3,$
and using ReLU activations. We train for 20 epochs using the Adam optimizer with learning rate $10^{-4}$ and batch size 64.
D.3.2 Data assimilation setup
To perform data assimilation with the L63 model, we generate scalar observations at each time step
$y_{n}=Hx_{n}+\eta_{n},\qquad H=\begin{bmatrix}1&0&0\end{bmatrix},\qquad\eta_{n}\sim\mathcal{N}(0,\sigma_{\mathrm{obs}}^{2}),\quad\sigma_{\mathrm{obs}}=5.$ (138)
At initialization, we sample $J$ particles from a Gaussian ball around the ground truth,
$x_{0}^{(j)}=x_{0}+\sigma_{\mathrm{init}}\,\xi^{(j)},\quad\xi^{(j)}\sim\mathcal{N}(\bm{0},\mathbf{I}),\quad\sigma_{\mathrm{init}}=5,\quad j=1,\ldots,J.$ (139)
D.3.3 Hyperparameter tuning
To tune the hyperparameters $t_{\min}$ and $\epsilon$ of DAISI, we use Bayesian optimization to minimize the CRPS on a separate trajectory, which we generate from the initial condition $x_{0}=(0,1,1.05)$ and integrate for $5000$ steps. We use DAISI with fixed values of $t_{\min}$ and $\epsilon$ and ensemble size $J=100$, to assimilate data on the last $200$ steps of the generated trajectory (this is to ensure the dynamics has reached statistical equilibrium), and evaluate its CRPS averaged across the last 100 steps (this is to ensure that the filter has stabilized). For Bayesian optimization, we used the following hyperpriors:
$\epsilon\sim\mathrm{LogUniform}\bigl(10^{-2},\,1\bigr),\qquad t_{\min}\sim\mathrm{Uniform}\bigl(10^{-2},\,0.99\bigr),$ (140)
and ran until convergence was observed.
D.3.4 Evaluation
After tuning the values for $t_{\min}$ and $\epsilon$, we evaluated DAISI on ten different trajectories starting from random initial conditions
$x_{0}^{(i)}\sim\mathcal{N}(\mu_{\text{train}},\sigma_{\text{train}}^{2}\mathbf{I}),\quad i=1,\ldots,10,$ (141)
and integrated for $5000$ time steps. We perform data assimilation with DAISI on the last $500$ steps of the generated trajectories, and evaluated the RMSE, CRPS and SSR, averaged across the last 100 steps of the assimilation window. As with the previous example, we used the Monte Carlo strategy for guidance with $J=10,000$ particles. For the bootstrap particle filter (BPF) baseline, we used $N_{p}=10,000$ particles to ensure high accuracy of the obtained results.
D.3.5 Ablation
In Figures 88, we plot the results of filtering using DAISI under different $(t_{\min},\epsilon)$-combinations. For reference, we also plot the result when no inverse sampling is performed in Figure 8 (i.e., it just samples from the posterior $\mathbb{P}_{\infty}^{{\bm{y}}}$ at every time step). This shows the importance of the inverse sampling step in DAISI to firstly, reduce the excessive sample variance when just using $\mathbb{P}_{\infty}^{{\bm{y}}}$ for assimilation, and secondly, for establishing time-continuity of the filter.
In Figure 8, we plot the results of DAISI with the default setting $t_{\min}=0.01$ and $\epsilon=0$. We see that while the filter is able to roughly track the overall pattern of the ground truth, the result is not very accurate, with narrow uncertainty bars that do not cover the ground truth consistently. This demonstrates the importance of tuning the hyperparameters $t_{\min}$ and $\epsilon$ to obtain an accurate filter. For example, Figure 8 displays the result of when just $t_{\min}$ is tuned, with $\epsilon$ fixed to 0 and Figure 8 displays the result when both hyperparameters are tuned. Both results show higher accuracy than the filter in Figure 8, highlighting the importance of tuning $t_{\min}$. Furthermore, the result in Figure 8 is more accurate than that of Figure 8, which shows how tuning $\epsilon$ can lead to further improvements in performance. Finally, in Figure 8, we display the results where we use the same value of $t_{\min}$ as used in Figure 8 (see Table 4), but setting $\epsilon$ to 0 to better observe the effect of $\epsilon$ on the filter. We clearly see that when $\epsilon$ is set to 0, the uncertainty bars become visibly narrower, making the predictions overconfident.
(a) No inversion
(b) No tuning
(c) Tuned $t_{\min}$ only
(d) Tuned both
(e) Tuned both, set $\epsilon=0$
Figure 8: Filtering results on the L63 system with DAISI under various hyperparameter settings.
$t_{\min}$ $\epsilon$
Figure 8 0.75
Figure 8 0.65 0.15
Table 4: Tuned hyperparameters used in the L63 experiments.
D.4 Surface Quasi-Geostrophic (SQG)
For the SQG experiments, we use the two-layer surface quasi-geostrophic model presented in (tulloch2009quasigeostrophic). This formulation evolves potential temperature at both the lower boundary ($z=0$) and an upper lid ($z=H$), capturing the influence of surface and tropopause-level dynamics on the interior flow. The three-dimensional geostrophic velocity field is diagnosed via a nonlocal spectral inversion, which couples the boundaries through stratified Green’s functions. Compared to classical SQG, which assumes decay into the deep interior ($z\to\infty$), the two-layer model confines dynamics to a finite slab and supports richer vertical structure, including baroclinic interactions.
The equations are solved numerically by first applying a fast Fourier transform (FFT) to map model variables to spectral space. They are then integrated forward with a fourth order Runge-Kutta solver that uses a 2/3 dealiasing rule and implicit treatment of hyperdiffusion. For more details, we refer to the GitHub repository of the model (sqgturb).
D.4.1 Model training
For the $64\times 64$ experiments, we generate 2000 training trajectories of 100 steps each at 3-hour intervals. Evaluation is done on 10 trajectories of 100 steps, and metrics are averaged over the final 20 steps. For the $256\times 256$ experiment, we generate 10 trajectories of length 1000 steps, but due to computational cost, evaluation is performed on a single 200-step run. All trajectories are initialized from approximate stationarity by spinning up the model for 300 days from random initial conditions. The data already has mean zero, but we normalize it with the standard deviation $\sigma_{\text{train}}=2660$.
Our backbone model for learning the drift $\bm{b}$} is a modified U-Net (UNET_Original) following the design of song2020score and Karras2022edm with 3.5M parameters. Since the SQG data is 2D periodic, we use circular padding to preserve spatial continuity. The same network configuration is used for both the $64\times 64$ and $256\times 256$ experiments, but each model is trained independently. The U-Net employs a hidden dimension of 32 throughout and consists of three hierarchical levels, with attention applied at the second level. Although we did not explore more advanced architectures, our framework is compatible with any model capable of learning a flow-matching objective.
The training process is executed in Pytorch, with setup and parameters detailed in Table 5.
Table 5: Optimizer Hyperparameters.
Optimizer hyperparameters
Optimiser AdamW (loshchilov_fixing_2017)
Initialization Xavier Uniform (glorot_understanding_2010)
LR decay schedule Cosine (loshchilov_sgdr_2017)
Peak LR 1e-3
Weight decay 1e-4
Warmup steps 1e3
Epochs 50
Batch size 200 (SQG 64), 2 (SQG 256), 10 (SEVIR)
D.4.2 Data assimilation setup
We perform data assimilation for all experiment setups in Table 2. The observations are given by
$y_{n}=\mathcal{H}(x_{n})+\eta_{n},\qquad\mathcal{H}(x_{n})=(A\circ h)(x_{n}),\qquad\eta_{n}\sim\mathcal{N}(0,\sigma_{\mathrm{obs}}^{2}),$ (142)
where $h$ is the operator in Table 2 and $A$ is the sparsity operator. The observation locations in $A$ are chosen uniformly at random before assimilation and remain the same across time and members.
Options for initializing DAISI.
There are several options for initializing DAISI, depending on the available initial conditions. In our numerical experiments, assimilation begins from a noisy version of the true initial state $\tilde{\bm{x}}_{0}\sim\mathcal{N}({\bm{x}}_{0},\sigma_{\text{init}})$. This perturbation is needed for LETKF to work, and is thus also used by DAISI to ensure fairness. In particular, this fairness is important for the squared observations, where the initial conditions affect how multimodal the filtering becomes.
Since these initial conditions will have unphysical noise, applying DAISI directly leads to unrealistic samples for the first few iterations. To address this, we perform an additional sampling step, running the forward SDE starting in the latent ${\bm{z}}_{t^{*}}=\alpha_{t^{*}}\tilde{\bm{x}}_{0}$, where $t^{*}\in[0,1]$ is chosen such that $\beta_{t^{*}}/\alpha_{t^{*}}=\sigma_{\text{init}}$. Conceptually, this treats the noisy initial condition as a partially inverted sample and lets DAISI remove the noise. We remark that this is an experimental detail and not a part of DAISI. If ones knowledge of the initial condition was an observation ${\bm{y}}_{0}$, one could start the assimilation by conditionally sampling from $\mathbb{P}_{\infty}^{{\bm{y}}_{0}}$.
This initialization procedure is reminiscent of SDEdit (meng2022sdedit), in which a sample is partially inverted by the stochastic interpolant ${\bm{z}}_{t}=\alpha_{t}{\bm{z}}_{1}+\beta_{t}{\bm{z}}_{0}$. This has been used for image editing, and in Table 11, we perform an ablation where it replaces the backward SDE in the inversion step.
D.4.3 Baselines
We compare DAISI against both classical and ML-based DA methods. Classical baselines include the Localized Ensemble Transport Filter (LETKF), while ML baselines include FlowDAS and the Ensemble Score Filter (EnSF).
EnSF:
We used the Ensemble Score Filter (EnSF) as described in (liang2025ensemble) with $\epsilon_{\alpha}=0.05$ and 1000 Euler steps. We note that the EnSF implementation includes an inflation step not mentioned in the original article, which we found crucial to prevent mode collapse. After each assimlation cycle, the particles are updated as
$\tilde{{\bm{x}}}^{(j)}_{n}=\bar{{\bm{x}}}_{n}+\frac{\sigma_{\text{init}}}{\sigma_{{{\bm{x}}}_{n}}}({{\bm{x}}}^{(j)}_{n}-\bar{{\bm{x}}}_{n}),$ (143)
where $\bar{{\bm{x}}}_{n}$ is the ensemble mean, $\sigma_{\text{init}}$ is the initial ensemble standard deviation and $\sigma_{{{\bm{x}}}_{n}}$ is the current ensemble standard deviation, averaged over all grid points.
FlowDAS:
We trained a conditional stochastic interpolant model for probabilistic forecasting (chen2024probabilistic), as used in FlowDAS, to emulate the forward dynamics of SQG at $64\times 64$ resolution. Our model conditions on the past six states for roll-out, i.e., generates samples from $p({\bm{x}}_{n}|{\bm{x}}_{n-1},\ldots,{\bm{x}}_{n-6})$. For the drift model $\bm{b}$}, we used a U-Net, similar to the architecture used for the interpolant in DAISI, with $64$ channels, and consisting of three hierarchical levels with channel multipliers $(1,2,2)$. Each level uses group-normalized residual blocks with 8 groups, and time embeddings are provided through learned sinusoidal features of dimension 32. Linear self-attention is applied at every resolution, together with a full multi-head self-attention block applied at the bottleneck between the encoder and decoder. The model is trained using AdamW with learning rate $2\times 10^{-4}$, cosine learning-rate decay, gradient-norm clipping, batch size 32, and dataset normalization with standard deviation $\sigma_{\text{train}}=2660$. Training is performed for 4096 epochs on 3-hourly SQG data.
For sampling, we use the parameters in Table 6, which are based on (chen2025flowdas) and finetuned for the different cases.
Table 6: Hyperparameter configurations for FlowDAS.
Experiment Guidance Strength $\zeta$ Euler steps MC guidance members $J$
Noisy 1 500 25
Sparse 3 500 25
Multimodal 0.5 500 25
Saturating 5 500 25
SEVIR 0.1 500 25
LETKF:
LETKF is a state-of-the-art DA method commonly used in the geosciences (wang2021multiscale). The hyperparameters used for LETKF are shown in Table 7. These are roughly based on the ones in (liang2025ensemble) but finetuned for the different cases. For the high-dimensional case, no parameter search is done due to computational cost.
Table 7: Hyperparameter configurations for LETKF.
Experiment Localization scale Relaxation to prior spread
Noisy $2500\text{\,}\mathrm{km}$ 0.5
Noisy 12h $1500\text{\,}\mathrm{km}$ 0.4
Sparse $1500\text{\,}\mathrm{km}$ 0.5
Multimodal $1500\text{\,}\mathrm{km}$ 0.5
Saturating $4500\text{\,}\mathrm{km}$ 0.9
High-dim. $1500\text{\,}\mathrm{km}$ 0.5
D.4.4 Hyperparameters
To identify suitable hyperparameters for DAISI, we conducted a greedy grid search across all experiments. The final hyperparameter settings for each experiment are listed in Table 8.
Table 8: Hyperparameter configurations for all SQG and SEVIR experiments.
Experiment Guide Method Solver Steps $\bm{\varepsilon}$ $\bm{t_{\min}}$ Guidance Strength $\zeta$
Noisy DPS 100 0.03 0.4 10
Noisy 12h MMPS 100 0.03 0.3 1
Sparse MMPS 100 0.03 0.3 1
Multimodal MMPS 100 0.03 0.3 1
Saturating MMPS 100 0.03 0.3 20
High-dim. MMPS 100 0.03 0.4 1
SEVIR MMPS 100 0.03 0.3 1
D.4.5 Results
We report RMSE and SSR for the SQG and SEVIR experiments. RMSE values in Table 9 align closely with the CRPS results in Table 3. The SSR in Table 10 indicates that DAISI is slightly overdispersive, but as shown in Figure 6, SSR is sensitive to the choice of $\epsilon$ and tuning this parameter more could further improve calibration.
Table 9: The RMSE for experiments on SQG and SEVIR. We display the mean and standard deviation across 10 independent trajectories, averaged over the last 20 (10 for SEVIR) steps. The best score for each experiment is highlighted in bold and the second best with an underline.
Experiment DAISI LETKF FlowDAS EnSF
Noisy $\mathbf{2.46}{\scriptstyle\pm 0.23}$ $\underline{2.51}{\scriptstyle\pm 0.26}$ $5.01{\scriptstyle\pm 0.69}$ $8.07{\scriptstyle\pm 1.01}$
Sparse $\mathbf{3.40}{\scriptstyle\pm 0.36}$ $\underline{4.55}{\scriptstyle\pm 0.53}$ $6.29{\scriptstyle\pm 0.90}$ $7.09{\scriptstyle\pm 0.74}$
Multimodal $\underline{3.71}{\scriptstyle\pm 0.95}$ $\mathbf{3.62}{\scriptstyle\pm 1.12}$ $6.93{\scriptstyle\pm 0.64}$ $9.64{\scriptstyle\pm 0.89}$
Saturating $\underline{2.96}{\scriptstyle\pm 0.17}$ $9.65{\scriptstyle\pm 0.69}$ $4.71{\scriptstyle\pm 0.69}$ $\mathbf{2.67}{\scriptstyle\pm 0.35}$
SEVIR $\mathbf{0.049}{\scriptstyle\pm 0.02}$ - $\underline{0.082}{\scriptstyle\pm 0.02}$ -
Table 10: The SSR for experiments on SQG and SEVIR. We display the mean and standard deviation across 10 independent trajectories, averaged over the last 20 (10 for SEVIR) steps. The best score for each experiment is highlighted in bold and the second best with an underline.
Experiment DAISI LETKF FlowDAS EnSF
Noisy $\textbf{1.06}{\scriptstyle\pm 0.03}$ $\underline{1.21}{\scriptstyle\pm 0.05}$ $0.78{\scriptstyle\pm 0.07}$ $0.49{\scriptstyle\pm 0.05}$
Sparse $1.24{\scriptstyle\pm 0.04}$ $\underline{1.23}{\scriptstyle\pm 0.10}$ $\textbf{1.13}{\scriptstyle\pm 0.06}$ $0.57{\scriptstyle\pm 0.05}$
Multimodal $1.66{\scriptstyle\pm 0.12}$ $\underline{0.76}{\scriptstyle\pm 0.15}$ $\textbf{1.21}{\scriptstyle\pm 0.07}$ $0.42{\scriptstyle\pm 0.04}$
Saturating $1.17{\scriptstyle\pm 0.03}$ $\underline{0.96}{\scriptstyle\pm 0.07}$ $\textbf{1.02}{\scriptstyle\pm 0.11}$ $1.60{\scriptstyle\pm 0.18}$
SEVIR $\underline{1.142}{\scriptstyle\pm 0.10}$ - $\textbf{1.085}{\scriptstyle\pm 0.13}$ -
D.4.6 Ablation
We assess the contribution of each component in DAISI through a series of ablations summarized in Table 11.
FlowDAS forecast:
Replacing the numerical model with the learned FlowDAS model yields almost identical performance to DAISI, indicating that our method is robust to this choice.
SDEdit backward step:
One can replace the inversion step with a single SDEdit (meng2022sdedit) step, in which the forecast ${\bm{z}}_{1}$ is partially noised to the latent ${\bm{z}}_{t_{\min}}=\alpha_{t_{\min}}{\bm{z}}_{1}+\beta_{t_{\min}}{\bm{z}}_{0}$, ${\bm{z}}_{0}\sim\mathcal{N}(\bm{0},\mathbf{I})$. Using a tuned value of $t_{\min}=0.4$, this approach performs worse than using the backward SDE. The reason is that $t_{\min}$ simultaneously controls the strength of the corrective guidance and how much of the forecast information is preserved. This coupling introduces a trade-off, starting later preserves more forecast information but leaves less room for correction.
No inversion:
Removing the inversion entirely results in a large degradation, underscoring the importance of injecting dynamical information from the forecast into the latent space. As illustrated in Figure 9, the inversion step is also critical for producing temporally smooth, physically consistent trajectories.
Hyperparameters:
We further study the sensitivity to the hyperparameters $\epsilon$ and $t_{\min}$ in the Sparse experiment. DAISI is generally robust to the choice of $\epsilon$, however, setting $\epsilon>0$ improves the probabilistic metrics (Figure 10). In contrast, $t_{\min}$ has a substantial impact on performance (Figure 11), suggesting its importance for minimizing the bias inherent in DAISI.
Assimilation interval:
Finally, we examine the effect of reduced assimilation frequency. Increasing the interval between observations amplifies forecast nonlinearities, which poses difficulties for LETKF. In the Noisy experiment, both LETKF and DAISI perform similarly with 3-hour assimilation intervals, but when the interval is extended to 12 hours, DAISI clearly outperforms LETKF (Figure 12).
Table 11: The CRPS for ablations on SQG. We display the mean and standard deviation across 10 independent trajectories, averaged over the last 20 (10 for SEVIR) steps.
Experiment DAISI with FlowDAS forecast with SDEdit backward step without Inversion
Noisy $1.32{\scriptstyle\pm 0.13}$ $1.34{\scriptstyle\pm 0.13}$ $2.12{\scriptstyle\pm 0.27}$ $2.63{\scriptstyle\pm 0.26}$
Sparse $1.73{\scriptstyle\pm 0.18}$ $1.72{\scriptstyle\pm 0.16}$ $2.17{\scriptstyle\pm 0.25}$ $2.40{\scriptstyle\pm 0.22}$
Multimodal $1.81{\scriptstyle\pm 0.44}$ $1.78{\scriptstyle\pm 0.41}$ $2.72{\scriptstyle\pm 0.45}$ $4.35{\scriptstyle\pm 0.33}$
Saturating $1.54{\scriptstyle\pm 0.09}$ $1.53{\scriptstyle\pm 0.11}$ $1.85{\scriptstyle\pm 0.20}$ $2.17{\scriptstyle\pm 0.19}$
Figure 9: Comparison of an ensemble member trajectory on the Noisy experiment between DAISI and DAISI* (without inversion). The DAISI* variant exhibits discontinuities between time steps t1t-1 and $t$, as forecast information is not propagated forward, leading to worse reconstructions and incoherent temporal evolution.
(a) The RMSE of the ensemble mean of the filtering distribution at each time step.
(b) The RMSE for each ensemble member at each time step.
(c) The spectral distance of each ensemble member compared to the ground truth and then averaged over the ensemble members.
(d) The CRPS of the empirical filtering distribution at each time step.
(e) The spread of the ensemble at each time step.
(f) The spread skill ratio of the ensemble at each time step.
Figure 10: The results for the $\epsilon$ ablation.
(a) The RMSE of the ensemble mean of the filtering distribution at each time step.
(b) The RMSE for each ensemble member at each time step.
(c) The spectral distance of each ensemble member compared to the ground truth and then averaged over the ensemble members.
(d) The CRPS of the empirical filtering distribution at each time step.
(e) The spread of the ensemble at each time step.
(f) The spread skill ratio of the ensemble at each time step.
Figure 11: The results for the $t_{\min}$ ablation.
(a) The RMSE of the ensemble mean of the filtering distribution at each time step.
(b) The RMSE for each ensemble member at each time step.
(c) The spectral distance of each ensemble member compared to the ground truth and then averaged over the ensemble members.
(d) The CRPS of the empirical filtering distribution at each time step.
(e) The spread of the ensemble at each time step.
(f) The spread skill ratio of the ensemble at each time step.
(g) The energy spectra at time step 1.
(h) The energy spectra at time step 50.
(i) The energy spectra at time step 100.
Figure 12: The results for the Noisy 12h experiment.
D.5 SEVIR
Here, we consider a real-life large-scale weather forecasting task using the Storm EVent Imagery and Radar (SEVIR) dataset (veillette2020sevir), which provides observations of severe convective storms across the continental United States. This specific experiment focuses on the Vertically Integrated Liquid (VIL) product, which serves as a 2-D proxy for precipitation intensity. Each data sample is a $128\times 128$ grid covering a $384\times$$384\text{\,}\mathrm{km}$ area at $2\text{\,}\mathrm{km}$ resolution, with snapshots recorded every $10\text{\,}\mathrm{min}$, for a total of $250\text{\,}\mathrm{min}$ (i.e., $25$ snapshots per sample).
For the forecast model, we used the checkpoint available in the official FlowDAS GitHub repository https://github.com/umjiayx/FlowDAS. This is based on a stochastic interpolant-based generative SDE for probabilistic forecasting, proposed in (chen2024probabilistic), with a U-Net backbone used for the drift. The model makes predictions autoregressively with inputs given by the past $6$ timesteps. For the guided FlowDAS baseline, we take the exact setup as in (chen2025flowdas), with parameters listed in Table 6. When beginning the assimilation, we use the first six ground truth states as initial conditions.
For DAISI we use the same U-Net as for the SQG experiments, without the circular padding. We do not scale the data as it is already in the range $[0,1]$.
Appendix E Additional figures
In Figures 1418 and 1924 we show additional samples and metrics for one of the assimilated trajectories.
Figure 13: A comparison of the ensemble mean, individual members, and ensemble standard deviation for each method at the last step of the assimilated trajectory for the Noisy experiment.
Figure 14: A comparison of the ensemble mean, individual members, and ensemble standard deviation for each method at the last step of the assimilated trajectory for the Sparse experiment.
Figure 15: A comparison of the ensemble mean, individual members, and ensemble standard deviation for each method at the last step of the assimilated trajectory of the Multimodal experiment.
Figure 16: A comparison of the ensemble mean, individual members, and ensemble standard deviation for each method at the last step of the assimilated trajectory of the Saturating experiment.
Figure 17: A comparison of the ensemble mean, individual members, and ensemble standard deviation for each method at the last step of the assimilated trajectory of the High-dim. experiment. The left colorbar is valid for the truth, samples and observation.
Figure 18: A comparison of the ensemble mean, individual members, and ensemble standard deviation for each method at the last step of the assimilated trajectory of the SEVIR experiment. The left colorbar is valid for the truth, samples and observation.
(a) The RMSE of the ensemble mean of the filtering distribution at each time step.
(b) The RMSE for each ensemble member at each time step.
(c) The spectral distance of each ensemble member compared to the ground truth and then averaged over the ensemble members.
(d) The CRPS of the empirical filtering distribution at each time step.
(e) The spread of the ensemble at each time step.
(f) The spread skill ratio of the ensemble at each time step.
(g) The energy spectra at time step 1.
(h) The energy spectra at time step 50.
(i) The energy spectra at time step 100.
Figure 19: The results for the Noisy experiment.
(a) The RMSE of the ensemble mean of the filtering distribution at each time step.
(b) The RMSE for each ensemble member at each time step.
(c) The spectral distance of each ensemble member compared to the ground truth and then averaged over the ensemble members.
(d) The CRPS of the empirical filtering distribution at each time step.
(e) The spread of the ensemble at each time step.
(f) The spread skill ratio of the ensemble at each time step.
(g) The energy spectra at time step 1.
(h) The energy spectra at time step 50.
(i) The energy spectra at time step 100.
Figure 20: The results for the Sparse experiment.
(a) The RMSE of the ensemble mean of the filtering distribution at each time step.
(b) The RMSE for each ensemble member at each time step.
(c) The spectral distance of each ensemble member compared to the ground truth and then averaged over the ensemble members.
(d) The CRPS of the empirical filtering distribution at each time step.
(e) The spread of the ensemble at each time step.
(f) The spread skill ratio of the ensemble at each time step.
(g) The energy spectra at time step 1.
(h) The energy spectra at time step 50.
(i) The energy spectra at time step 100.
Figure 21: The results for the Multimodal experiment.
(a) The RMSE of the ensemble mean of the filtering distribution at each time step.
(b) The RMSE for each ensemble member at each time step.
(c) The spectral distance of each ensemble member compared to the ground truth and then averaged over the ensemble members.
(d) The CRPS of the empirical filtering distribution at each time step.
(e) The spread of the ensemble at each time step.
(f) The spread skill ratio of the ensemble at each time step.
(g) The energy spectra at time step 1.
(h) The energy spectra at time step 50.
(i) The energy spectra at time step 100.
Figure 22: The results for the Saturating experiment.
(a) The RMSE of the ensemble mean of the filtering distribution at each time step.
(b) The RMSE for each ensemble member at each time step.
(c) The spectral distance of each ensemble member compared to the ground truth and then averaged over the ensemble members.
(d) The CRPS of the empirical filtering distribution at each time step.
(e) The spread of the ensemble at each time step.
(f) The spread skill ratio of the ensemble at each time step.
(g) The energy spectra at time step 1.
(h) The energy spectra at time step 100.
(i) The energy spectra at time step 200.
Figure 23: The results for the High-dim. experiment.
(a) The RMSE of the ensemble mean of the filtering distribution at each time step.
(b) The RMSE for each ensemble member at each time step.
(c) The spectral distance of each ensemble member compared to the ground truth and then averaged over the ensemble members.
(d) The CRPS of the empirical filtering distribution at each time step.
(e) The spread of the ensemble at each time step.
(f) The spread skill ratio of the ensemble at each time step.
(g) The energy spectra at time step 1.
(h) The energy spectra at time step 10.
(i) The energy spectra at time step 19.
Figure 24: The results for the SEVIR experiment.